### 13 Mar 2026 v1

# Collaborative Multi-Agent Optimization for Personalized Memory System

Abstract
Memory systems are crucial to personalized LLMs by mitigating the context window limitation in capturing long-term user-LLM conversations. Typically, such systems leverage multiple agents to handle multi-granular memory construction and personalized memory retrieval tasks. To optimize the system, existing methods focus on specializing agents on their local tasks independently via prompt engineering or fine-tuning. However, they overlook cross-agent collaboration, where independent optimization on local agents hardly guarantees the global system performance. To address this issue, we propose a Collaborative Reinforcement Learning Framework for Multi-Agent Memory Systems (CoMAM), jointly optimizing local agents to facilitate collaboration. Specifically, we regularize agents’ execution as a sequential Markov decision process (MDP) to embed inter-agent dependencies into the state transition, yielding both local task rewards (e.g., information coverage for memory construction) and global rewards (i.e., query-answer accuracy). Then, we quantify each agent’s contribution via group-level ranking consistency between local and global rewards, treating them as adaptive weights to assign global credit and integrate local-global rewards. Each agent is optimized by these integrated rewards, aligning local improvements with the global performance. Experiments show CoMAM outperforms leading memory systems, validating the efficacy of our proposed collaborative reinforcement learning for joint optimization.
Multi-agent, Memory System, Reinforcement Learning
1 Introduction
Figure 1: Construction and retrieval agents, optimized on local tasks independently, yield lower global system performance than those under joint optimization.
Memory systems (Chhikara et al., 2025; Zhong et al., 2024; Hu et al., 2025) have emerged as a pivotal frontier for personalized large language models (LLMs), which preserve users’ long-term conversational history beyond LLMs’ context window and selectively integrate it to tailor personalized responses (Chen et al., 2024a; Zhang et al., 2025). Typically, the memory system comprises multiple specialized agents: several construction agents summarize and store user history into multi-granular memories (e.g., long/short or coarse/fine), respectively (Li et al., 2025; Wang & Chen, 2025; Wang et al., 2025); and retrieval agents select and integrate relevant memories for subsequent queries via RAG (Lewis et al., 2020) or search-based reasoning (Yu et al., 2026; Jin et al., 2025). To enhance system effectiveness, existing methods focus on optimizing these agents for their respective tasks independently: Prompt engineering approaches (Li et al., 2025; Zhong et al., 2024) specialize each agent for its local task with well-crafted prompts, while reinforcement learning methods (Wang et al., 2025; Zhou et al., 2026; Yan et al., 2025b) train agents separately using rewards such as the query-answer accuracy.
Despite local successes, they overlook cross-agent collaboration, treating the memory system as a “collection of local experts” rather than a collaborative team that evolves together for a global goal. Critically, independent optimization on local agents hardly guarantees global system performance (Akella, 2025; Wu et al., 2026). While individual agents may excel at their respective tasks, their independent strategies can conflict when combined in the global pipeline. To be more specific, if the fine-grained construction agent is optimized by treating retrieval agents as fixed environment components, it may store redundant or noisy content that impairs retrieval effectiveness. Conversely, if retrieval agents are trained without considering construction strategies, they may over-filter information and miss critical details during retrieval. As illustrated in Figure 1, even with independently well-trained agents (rewards converged), the combined system consistently underperforms that with joint optimization across different conversational history lengths. This motivates us to explore jointly optimizing multiple agents in the memory system to facilitate collaboration, which faces two challenges (see Figure 2):
Figure 2: Illustration of challenges for joint optimization.
  • Heterogeneity and Asynchrony: Multi-granular construction and retrieval agents operate with heterogeneous configurations and asynchronous executions across different tasks. Their discrete parameter space and inherent structural complexity complicate end-to-end optimization.
  • Local-Global Alignment: Fostering collaboration requires aligning local improvement with the global performance. However, directly assigning equal shares of the global objective to individual agents introduces ambiguity in task execution, as agents’ contributions are divergent.
To address these challenges, we introduce CoMAM, a Collaborative reinforcement learning framework for Multi-Agent Memory systems. Specifically, we formulate a standard memory system with three specialized agents for multi-granular memory construction and retrieval: an extraction agent to construct fine-grained memories (e.g., key events and facts), a profile agent to construct coarse-grained memories (e.g., users’ personalized preferences), and a retrieval agent to retrieve memories and answer queries. To optimize these heterogeneous agents jointly, we frame their asynchronous execution as a sequential Markov decision process (MDP), which embeds inter-agent dependencies into state transitions and yields both local task-specific rewards (e.g., information coverage for memory construction) and global system rewards (i.e., query answer accuracy). To achieve local-global alignment, CoMAM quantifies each agent’s system contribution by measuring group-level ranking consistency between its local and global rewards, then adaptively allocates global rewards to agents based on this consistency. Each agent is ultimately optimized with a combination of its task-specific reward and adaptively assigned global reward share, driving end-to-end collaboration toward the global objective. Extensive experiments on long-term personalized conversational tasks show CoMAM outperforms leading memory baselines, validating the efficacy of our collaborative reinforcement learning for jointly optimizing multi-agent memory systems.
2 Related work
Memory System for Personalized LLMs
[…]
Optimization of Multi-agent System
Multi-agent Systems (MAS) (Sun et al., 2024; Chen et al., 2024b) leverage multiple specialized agents with heterogeneous configurations and asynchronous execution to tackle complex tasks beyond the capacity of a single agent. To boost MAS performance, existing methods focus on independent optimization, primarily via prompt engineering (Patil et al., 2024; Wu et al., 2024) or fine-tuning of agents on specific tasks (Subramaniam et al., 2025; Yan et al., 2025b). Yet such methods yield only marginal gains, as they overlook the value of joint optimization for inter-agent collaboration. Consequently, recent research has shifted to reinforcement learning (RL)-based collaborative training for end-to-end system optimization, such as MAPoRL (Park et al., 2025) and MARFT (Liao et al., 2025).
Credit Assignment
Credit assignment(Zhou et al., 2020; Bo et al., 2024a; Tseng et al., 2022) concerns properly attributing a team’s global outcome to the specific agents in Multi-Agent Reinforcement Learning (MARL). This is particularly difficult in quantifying agents’ contributions to the outcome, which is often obscured by the actions of other agents and environmental noise. Naive credit assignment approaches evenly distribute global team rewards across agents (Zhao et al., 2026; Hong et al., 2025). Though simple, this leads to significant credit assignment ambiguity: agents cannot distinguish their specific impact, resulting in high variance in policy updates and unstable convergence.
3 Preliminary and Challenges
3.1 Task Formulation for Multi-agent Memory System
We define the user’s long-term conversational history as $\mathcal{H}=\{h_{1},\dots,h_{N}\}$, which consists of $N$-turn interactions that exceed the LLM’s context window. To facilitate personalization, the Memory System transforms the history $\mathcal{H}$ into multi-granular memories $\mathcal{M}$ that can be retrieved to answer users’ subsequent queries $q$, realized by diverse agents on construction and retrieval tasks (Hu et al., 2025):
  • Memory Construction: The agents (Li et al., 2025) under policies $\{\pi_{\theta}^{\text{cons}_{g}}\}$, distill the raw conversational history $\mathcal{H}$ at different granularities $g$, respectively, finally constructing a multi-granular memory set $M$:
    $\mathcal{M}_{g}=\pi_{\theta}^{\text{cons}_{g}}(\mathcal{H}),$ (1)
    $\mathcal{M}=\{\mathcal{M}_{g}\}.$ (2)
  • Memory Retrieval: Given a subsequent query $q$, retrieval agents (Zhou et al., 2026) under policies $\{\pi_{\theta}^{\text{ret}}\}$ identify the $K$ most relevant memories from $\mathcal{M}$. Then, the agent with policy $\pi_{\theta}^{\text{LLM}}$ generates a response $p$ for the query $q$ using the retrieved $K$ memories. Formally:
    $\{\bm{m}_{k}\}_{k=1}^{K}=\pi_{\theta}^{\text{ret}}(q,\mathcal{M},K),$ (3)
    $p=\pi_{\theta}^{\text{LLM}}\left(q,\{\bm{m}_{k}\}_{k=1}^{K}\right).$ (4)
    For efficiency, $\pi_{\theta}^{\text{ret}}$ and $\pi_{\theta}^{\text{LLM}}$ can be unified into a single policy that performs both memory retrieval and personalized response generation for query $q$.
The global goal of the memory system is to align the generated response $p$ with the user’s personalized preferences and experiences. Formally, given a dataset $\mathcal{D}$ with input pairs $(\mathcal{H},q)$ and a performance evaluation function $\mathcal{F}$, the optimization objective is to find the optimal policy set $\Pi^{*}$ that maximizes the expected global performance:
$\Pi^{*}=\arg\max_{\Pi}\mathbb{E}_{(\mathcal{H},q)\sim\mathcal{D}}\left[\mathcal{F}\left(p^{*},\Pi(\mathcal{H},q)\right)\right].$ (5)
Here, $\Pi=\{\pi_{\theta}^{\text{cons}_{g}},\pi_{\theta}^{\text{ret}}\}$ denotes the policy set, and $p^{*}$ denotes the ground-truth response.
3.2 Reinforcement Learning for LLMs
To optimize LLMs’ policy $\pi_{\theta}$, Reinforcement Learning (RL) (Ouyang et al., 2022; Zheng et al., 2023) is typically used to maximize the expected reward while maintaining stability relative to a reference model $\pi_{\text{ref}}$. In this work, we primarily build upon Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an efficient RL algorithm that eliminates the need for a separate value function (critic) through group-based relative rewards. Given the input $x$, GRPO samples a group of outputs $\left\{o_{1},\ldots,o_{G}\right\}$ from the current policy $\pi_{\theta}$ and optimizes the objective function:
$\mathcal{J}_{GRPO}\left(\pi_{\theta}\right)=$ (6)
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left\{\min\left[\rho_{i,t}\hat{A}_{i,t},\hat{\rho}_{i,t}\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}\right\},$
where $\rho_{i,t}=\frac{\pi_{\theta}\left(o_{i,t}\mid x,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid x,o_{i,<t}\right)}$ is the importance sampling ratio, $\hat{\rho}_{i,t}=\operatorname{clip}\left(\rho_{i,t};1-\epsilon,1+\epsilon\right)$ is the clipped ratio. The advantage $\hat{A}_{i,t}$ for each output $o_{i}$ is computed by normalizing its reward $r_{i}$ relative to the group mean and the standard deviation. Formally, we have:
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)}.$ (7)
The KL divergence is defined as:
$\mathbb{D}_{KL}=\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-1.$ (8)
4 Method
Figure 3: The overview of our proposed framework, CoMAM, which regularizes agents’ execution as MDP trajectories for joint RL optimization and fosters collaboration via adaptive credit assignment to achieve local-global alignment.
In this section, we present our proposed CoMAM, a Collaborative Reinforcement Learning Framework designed to address the key challenges of jointly optimizing the multi-agent memory systems, as shown in Figure 3. We begin by defining the task-specific agents alongside the local and global rewards in Sec 4.1. Then, we elaborate on regularizing the execution trajectory as a Markov decision process (MDP) in Sec 4.2, which enables joint optimization despite agent heterogeneity and asynchrony. Finally, we present the adaptive credit assignment mechanism in Sec 4.3, which quantifies individual agent contributions and aligns local improvements with global performance.
4.1 Task-Specific Agents Definition
Building on existing multi-agent memory paradigms (Hu et al., 2025; Li et al., 2025), we define a standard memory system consisting of three specialized agents: Extraction Agent, Profile Agent, and Retrieval Agent.
Extraction Agent extracts fine-grained memory $\mathcal{M}_{f}$ with policy $\pi_{\theta}^{\text{cons}_{f}}$ from the raw conversational history $\mathcal{H}$. Formally, we define the set of evident information relevant to query $q$ as $\mathcal{E}_{q}\subseteq\mathcal{H}$. The agent’s objective is to maximize relevant evidence coverage with the reward function:
$r_{\text{cons}_{f}}=\alpha\cdot\frac{|{\mathcal{M}_{f}}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\alpha)\cdot\frac{|{\mathcal{M}_{f}}\cap\mathcal{E}_{q}|}{|{\mathcal{M}_{f}}\cup\mathcal{E}_{q}|},$ (9)
where $r_{\text{cons}_{f}}$ balances the coverage of relevant information and the filtering of irrelevant content. To prioritize coverage, $\alpha$ is set to $0.8$.
Profile Agent abstracts coarse-grained memory $\mathcal{M}_{c}$ (e.g., user’s preferences or behavioral patterns) from $\mathcal{M}_{f}$, guided by policy $\pi_{\theta}^{\text{cons}_{c}}$. The agent aims to maximize abstraction rationality, quantified by the model-based reward function:
$r_{\text{cons}_{c}}=\mathcal{V}(\mathcal{M}_{c},\mathcal{H})\in[0,1],$ (10)
where $r_{\text{cons}_{c}}$ evaluates the quality of $\mathcal{M}_{c}$ relative to the conversation history $\mathcal{H}$ according to predefined rubrics (e.g., scenario, preference, and user profile). $\mathcal{V}$ denotes a frozen pre-trained LLM for reward generation. Finally, the comprehensive memory set is defined as $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$.
Retrieval Agent retrieves top-$k$ relevant memories $\mathcal{E}_{q}^{\prime}=\{\bm{m}_{k}\}_{k=1}^{K}$ from memory set $\mathcal{M}$ and generate personalized response $p$ for user’s subsequent query $q$, guided by policy $\pi_{\theta}^{\text{ret}}$. To evaluate the retrieval performance, we define a rule-based reward function that enables precise matching between the recalled $\mathcal{E}_{q}^{\prime}$ and ground-truth evidence $\mathcal{E}_{q}$:
$r_{\text{ret}}=\beta\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\beta)\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}^{\prime}\cup\mathcal{E}_{q}|},$ (11)
where $r_{\text{ret}}$ balances the coverage and precision of retrieval. To prioritize retrieval precision while minimizing noise, $\beta$ is set to $0.2$.
To evaluate the memory system’s global performance, we define the rule-based reward for $\pi_{\theta}^{\text{ret}}$ to evaluate its answering accuracy for query $q$. Formally, let $p^{*}$ denote the ground-truth response for the personalized query $q$, we have:
$r_{\text{ans}}=\mathbb{I}\left(p^{*}=p\right)$ (12)
where $\mathbb{I}$ denotes the indicator function that measures alignment between the generated response $p$ and ground-truth answer $p^{*}$, yielding $1$ if $p^{*}=p$ and 0 otherwise.
Notably, the number and tasks of these agents can be further expanded to meet diverse application requirements in a multi-agent memory system.
4.2 Execution Trajectory Regularization
Usually, construction agents run periodically to manage historical conversations, while retrieval agents are activated for incoming queries. To enable joint training under such heterogeneous and asynchronous settings, we model the entire construction and retrieval pipeline of the multi-agent memory system as a sequential Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},T,R)$, where each step is executed by a heterogeneous agent. Formally, let $\pi_{\theta}^{0}=\pi_{\theta}^{\text{cons}_{f}}$, $\pi_{\theta}^{1}=\pi_{\theta}^{\text{cons}_{c}}$$\pi_{\theta}^{2}=\pi_{\theta}^{\text{ret}}$ denote the policies at step $n=1,2,3$, respectively. Then, for each MDP trajectory $\tau$, we have:
  • $\mathcal{S}$: State space, capturing the system state at each step in the MDP. Specifically, $s_{0}=\mathcal{H}$ denotes the Extraction Agent’s initial state of raw conversational history $\mathcal{H}$. $s_{1}=\mathcal{M}_{f}$ denotes the Profile Agent’s state, where $\mathcal{M}_{f}$ is the fine-grained memory extracted by the prior Extraction Agent. $s_{2}=(\mathcal{M},q)$ denotes the Retrieval Agent’s state, with $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$ as the multi-granular memory set and $q$ as the incoming user query. $s_{3}=(\mathcal{E}_{q}^{\prime},p)$ denotes the system’s final state, comprising retrieved information $\mathcal{E}_{q}$ and generated response $p$.
  • $\mathcal{A}$: Action space, defining executable actions of each heterogeneous agent in the MDP. Specifically, $a_{0}$ denotes extracting fine-grained memory by the Extraction Agent. $a_{1}$ denotes abstracting coarse-grained memory by the Profile Agent. $a_{2}$ denotes the Retrieval Agent’s action, which involves retrieving top-$k$ relevant memories and generating a personalized response.
  • $T$: State transition function, where $s_{n+1}=T(s_{n},a_{n})$ and $a_{n}=\pi_{\theta}^{n}(s_{n})$. Specifically, the Extraction Agent transitions to the Profile Agent via $s_{1}=T(s_{0},a_{0})$ and $a_{0}=\pi_{\theta}^{0}(\mathcal{H})$; the Profile Agent then transitions to the Retrieval Agent through $s_{2}=T(s_{1},a_{1})$ and $a_{1}=\pi_{\theta}^{1}(\mathcal{M}_{f})$; finally, the Retrieval Agent produces the system’s final state $s_{3}=T(s_{2},a_{2})$ where $a_{2}=\pi_{\theta}^{2}(\mathcal{M},q)$.
  • $R$: Reward function, quantifying the rewards of agent actions in the MDP. As detailed in Sec 4.1, the local task rewards for each step are $r_{0}=r_{\text{cons}_{f}}$ for the Extraction Agent, $r_{1}=r_{\text{cons}_{c}}$ for the Profile Agent, and $r_{2}=r_{\text{ret}}$ for the Retrieval Agent. And, the global system reward is defined as $r_{3}=r_{\text{ans}}$.
Within this formulation, CoMAM embeds inter-agent dependencies into state transitions along the optimization trajectory. The multi-agent memory system operates sequentially to generate local and global rewards, enabling end-to-end RL optimization with simultaneous updates of heterogeneous policies. The task-specific capabilities and inter-agent dependencies learned in this MDP enable agents to perform well in more complex inference environments, where divergent operating frequencies demand enhanced collaboration.
4.3 Adaptive Credit Assignment Mechanism
To align local improvements with global performance during joint optimization, we propose an adaptive credit assignment mechanism to allocate global system rewards to local agents. Instead of assigning global rewards equally to each agent (Hong et al., 2025; Zhao et al., 2026), CoMAM quantifies agents’ varying contributions to the system and treats them as adaptive weights for allocation.
Formally, for each input $(\mathcal{H},q)$, we sample a group of MDP trajectories $\{\tau_{i}\}_{i=1}^{G}$, collecting the local rewards for each agent $n$, $\mathcal{R}_{\text{local},n}=\{r_{n}^{(i)}\}_{i=1}^{G}$, and the corresponding global rewards, $\mathcal{R}_{\text{global}}=\{r_{3}^{(i)}\}_{i=1}^{G}$, where $n\in\{0,1,2\}$.
To quantify agent $n$’s contribution to the memory system, we compute the ranking consistency score between $\mathcal{R}_{\text{local},n}$ and $\mathcal{R}_{\text{global}}$ using the Normalized Discounted Cumulative Gain (NDCG) metric. Specifically, the NDCG-based consistency score for agent $n$ is defined as:
$v_{n}=\text{NDCG}(\sigma(\mathcal{R}_{\text{local},n}),\sigma(\mathcal{R}_{\text{global}})),$ (13)
where the global reward ranking $\sigma(\mathcal{R}_{\text{global}})$ serves as the ideal ranking for the local reward ranking $\sigma(\mathcal{R}_{\text{local},n})$. We then normalize $v_{n}$ to obtain adaptive credit weights $w_{n}$ that satisfy $\sum_{n}w_{n}=1$:
$w_{n}=\frac{\exp(v_{n})}{\sum_{n^{\prime}}\exp(v_{n^{\prime}})}.$ (14)
The final optimization signal for agent $n$ is the combination of its local rewards and allocated global credit, weighted by the adaptive weight $w_{n}$:
$r_{\text{final},n}^{(i)}=r_{n}^{(i)}+w_{n}\cdot r_{3}^{(i)}.$ (15)
After the adaptive credit assignment, we apply the GRPO algorithm (as defined in Sec 3.2) on each agent based on the integrated rewards $r_{\text{final}}$. Formally, the objective function for the agent $n$ is defined as:
$\mathcal{J}_{GRPO}\left(\pi_{\theta}^{n}\right)=$ (16)
$\frac{1}{G}\sum_{i=1}^{G}\min\left[\rho_{n}^{(i)}A_{n}^{(i)},\;\text{clip}\left(\rho_{n}^{(i)},\;1\pm\epsilon\right)A_{n}^{(i)}\right],$
where $\rho_{n}^{(i)}=\frac{\pi_{\theta}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}{\pi_{\theta_{\text{old}}}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}$ is the importance sampling ratio, the advantage $A_{n}^{(i)}$ for each agent’s action $a_{n}^{(i)}$ is computed by normalizing its reward $r_{\text{final},n}^{(i)}$. By optimizing each agent with both local task-specific rewards and allocated global system rewards, we can align local improvements with global performance and foster collaboration.
5 Experiments
In this section, we conduct extensive experiments to validate the effectiveness of our proposed CoMAM by addressing the following questions:
  • RQ1: How does CoMAM perform against leading baselines of memory systems for personalization?
  • RQ2: What is the contribution of each agent within the memory system?
  • RQ3: What is the impact of regularizing agents’ execution as an MDP for jointly optimizing the memory system?
  • RQ4: What is the contribution of adaptive credit assignment for facilitating collaboration in the memory system?
  • RQ5: What is the computational cost of the CoMAM?
Table 1: Query-Answering Accuracy on the PersonaMem Benchmark across 32K, 128K, and 1M context length settings. Bold indicates the best performance, and underlining denotes the second best.
Method Qwen Llama
32K 128K 1M 32K 128K 32M
No memory
Base 0.41 0.39 0.38 0.35 0.31 0.32
RAG (Lewis et al., 2020) 0.48 0.45 0.41 0.43 0.39 0.36
Prompt-based memory
CAM (Li et al., 2025) 0.53 0.50 0.45 0.48 0.45 0.43
MemoryBank (Zhong et al., 2024) 0.51 0.47 0.44 0.45 0.42 0.39
A-Mem (Xu et al., 2025) 0.50 0.44 0.45 0.46 0.37 0.36
RL-based memory
Mem1 (Zhou et al., 2026) 0.59 0.57 0.58 0.56 0.58 0.61
Memory-R1 (Yan et al., 2025b) 0.58 0.60 0.60 0.57 0.61 0.60
Ours 0.64 0.70 0.66 0.62 0.68 0.69
Impro +8.5% +16.7% +10% +8.8% +11.5% +13.1%
5.1 Experimental Settings
Dataset
We conduct experiments on the PersonaMem benchmark (Jiang et al., 2025), a large-scale dataset including over 180 long-term user-LLM interaction histories and about 6,000 multiple-choice questions as subsequent queries. The benchmark is split into three experimental settings with varying history lengths (i.e., 32K, 128K, and 1M tokens), each paired with subsequent queries spanning seven distinct types: (1) Recall user-shared facts, (2) Suggest new ideas, (3) Acknowledge latest user preferences, (4) Track full preference evolution, (5) Revisit reasons for preference updates, (6) Provide preference-aligned recommendations, and (7) Generalize to new scenarios. Compared to existing benchmarks (Maharana et al., 2024; Wu et al., 2025), PersonaMem’s strengths in long context history and personalized query type enable rigorous assessment of memory systems in personalized conversational scenarios. Dataset processing details are provided in Appendix B.1.
Baselines
We conduct a comprehensive evaluation of CoMAM against multiple leading baselines, including methods without a memory system “No memory”, methods of building memory systems via dedicated prompts “Prompt-based memory”, and methods of optimizing the memory systems with RL “RL-based memory”. Detailed introductions and implementations of baselines are provided in Appendix B.
  • No memory: “Base” represents directly injecting raw conversations as LLMs’ input for query answering; RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history;
  • Prompt-based memory: A-Mem (Xu et al., 2025) instantiates a memory system with construction, evolution, and retrieval agents; CAM (Li et al., 2025) and MemoryBank (Zhong et al., 2024) build multi-level (fine-to-coarse granularity) memory via multi-agent dedicated prompts.
  • RL-based memory: Mem1 (Zhou et al., 2026) uses RL to optimize agents for memory searching during reasoning; Memory-R1 (Yan et al., 2025b) separately optimizes memory construction and retriever agents via RL.
Figure 4: Detailed performance of different methods across seven question types on the PersonaMem benchmark across three history length settings (i.e., 32K, 128K, and 1M). The number of question types is consistent with those listed in Section 5.1
Implementation Details.
To evaluate CoMAM’s effectiveness and generalizability, we conduct experiments on PersonaMem across its three history length settings (i.e., 32K, 128K, and 1M tokens) using LLM backbones from the Qwen (Bai et al., 2023) and Llama (Touvron et al., 2023) families. Specifically, the extraction agent, profile agent, and reward model $\mathcal{V}$ are instantiated with Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct, while the retrieval agent leverages Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct, with model sizes tailored to each agent’s task complexity. All LLM backbones are configured with a 32K-token context window, and we adopt the average query-answer accuracy as the personalization evaluation metric. Table 1 presents the main results for both Qwen and Llama; other tables show only Qwen for simplicity. All tabulated results are averages of at least three independent runs with distinct random seeds. More details are provided in the Appendix B.3.
5.2 Overall Performance (RQ1)
To validate CoMAM’s efficacy for personalized LLMs, we compare its query-answer accuracy against leading memory system baselines on the PersonaMem benchmark. As shown in Table 1, CoMAM delivers superior performance across all context length configurations (32K, 128K, and 1M) and two LLM backbone families (Qwen and Llama), fully confirming CoMAM’s effectiveness and generalizability in personalized conversational scenarios. Notably, RL-based memory baselines outperform prompt-based memory systems, validating that reinforcement learning is effective for optimizing memory system agents. For instance, the top prompt-based baseline CAM (Li et al., 2025) achieves only 0.450.45-$0.53$ accuracy across contexts, while the RL-based Memory-R1 (Yan et al., 2025b) maintains a higher accuracy range of 0.570.570.610.61. Moreover, CoMAM outperforms the state-of-the-art baseline Memory-R1, which employs independent agent optimization. This highlights the critical importance of joint optimization for multi-agent memory systems to achieve collaboration. Figure 4 details Qwen-based results for the seven question types, and we observe that CoMAM performs well in both detailed factual recall (e.g., Type 1) and abstract preference-based question types (e.g., Type 6), providing additional validation of its efficacy.
Table 2: Ablation study for the contribution of different agents to the memory system.
   Method    Accuracy (%)
   32K    128K    1M
   Base    0.23    0.26    0.37
   w/o Ex    0.52    0.52    0.51
   w/o Pr    0.57    0.52    0.57
   w/o Re    0.54    0.51    0.51
   mo train    0.50    0.46    0.44
   w/o train Ex    0.57    0.61    0.54
   w/o train Pr    0.57    0.61    0.64
   w/o train Re    0.62    0.66    0.59
   Ours    0.64    0.70    0.66
Table 3: Ablation study of MDP regularization for heterogeneous, asynchronous agents. The “no train” variant denotes a memory system composed of agents merely guided by prompting.
Method Accuracy (%)
32K 128K 1M
single policy
no train 0.51 0.48 0.45
+ independent RL 0.57 0.56 0.54
+ MDP RL 0.60 0.59 0.59
heterogeneous policy
no train 0.50 0.46 0.44
+ independent RL 0.57 0.61 0.59
+ MDP RL (ours) 0.64 0.70 0.66
5.3 Ablation for Agents (RQ2)
To verify the necessity and contribution of the three core agents (Extraction Agent, Profile Agent, and Retrieval Agent), we designed targeted ablation experiments, with results presented in Table 2. First, the “Base” baseline (without memory agents) achieves the lowest performance across all context lengths, confirming the significance of memory systems for long-context personalized conversations. Second, removing any agent can lead to significant performance degradation, whether the Extraction Agent (“w/o Ex”) that builds factual fine-grained memories, the Profile Agent (“w/o Pr”) that captures core user preferences, or the Retrieval Agent (“w/o Re”) that searches for memories to answer queries. These results demonstrate that all three agents play complementary and unique roles in supporting the personalized memory system. Furthermore, the “w/o train” variants (full agent architecture with one agent left untrained) provide additional validation: untraining any single agent results in 2–8% performance drops compared to CoMAM, highlighting that optimizing these agents is critical to building an effective memory system.
5.4 Ablation for MDP Regularization (RQ3)
To address heterogeneity and asynchrony for joint optimization, we regularize agents’ execution as a sequential MDP trajectory. To validate MDP Regularization’s efficacy, we design “single policy” (all agents share one policy across tasks (Yu et al., 2026; Bo et al., 2024b)) and “heterogeneous policy” (each agent has a dedicated policy per task) variants, with two training strategies: “independent RL” (fine-tunes one agent’s policy at a time, others fixed without updates) and “MDP RL” (simultaneously updates policies via MDP trajectory for joint optimization). As shown in Table 3, “single policy” variants consistently underperform “heterogeneous policy” variants, confirming that a shared policy undermines agent specialization by integrating diverse task-specific objectives. Furthermore, “+MDP RL” outperforms “independent RL” across both policy paradigms, validating that MDP regularization explicitly models inter-agent sequential dependencies and enables both local task specialization and global collaboration.
5.5 Ablation for Adaptive Credit Assignment (RQ4)
To address RQ4, we design diverse credit assignment strategies to validate CoMAM’s adaptive mechanism. As shown in Table 4, optimizing agents solely with local task-specific rewards (“MDP RL w/ L”) or global rewards (“MDP RL w/ G”) yields performance improvements over the “no train” variant and “independent RL” variant, but lags behind variants that integrate global rewards with local rewards (“MDP RL w/ LG” and “Ours”). This demonstrates the necessity of combining local task rewards and global rewards to balance agent specialization on local tasks and collaboration for global performance. Notably, “Ours” (which is equipped with adaptive credit assignment) outperforms “MDP RL w/ LG” (which allocates equal shares of global rewards to local tasks), confirming the critical value of adaptively aligning local improvements with global performance. As illustrated in Figure 5, our adaptive weight $w_{n}$ for credit assignment outperforms any fixed weights, further validating the efficacy of the adaptive mechanism. Moreover, adaptive credit assignment can also enhance agents’ task-specific ability: our method outperforms both the “w/ LG” and “w/ L” on local task performance, validating this effect.
Table 4: Ablation study for different credit assignment strategies.
Method Accuracy (%)
32K 128K 1M
no train 0.50 0.46 0.44
+ independent RL 0.57 0.61 0.59
MDP RL w/ L 0.57 0.65 0.61
MDP RL w/ G 0.60 0.64 0.60
MDP RL w/ LG 0.62 0.65 0.63
Ours 0.64 0.70 0.66
5.6 Efficiency (RQ5)
We further evaluate the training efficiency of CoMAM by comparing the convergence steps of joint optimization and independent optimization across 32K, 128K, and 1M context length settings, with results presented in Table 5. The total convergence steps serve as the core metric to reflect training efficiency, where fewer steps indicate higher efficiency in reaching stable performance. Specifically, independent optimization trains each agent separately, with total steps being the sum of individual convergence steps for the Extraction, Profile, and Retrieval Agents. In contrast, joint optimization trains all agents in parallel, with total steps determined by the slowest-converging agent, resulting in fewer total steps. While individual agents under joint training may take more steps (e.g., 58 vs. 48 steps for the Extraction Agent at the 32K setting), parallel training shortens the total training duration. Notably, inference latency is decoupled from the training paradigm: agents optimized by joint or independent optimization achieve comparable inference efficiency at deployment. Yet joint optimization delivers superior global performance via inter-agent collaboration, highlighting the distinct advantage of joint multi-agent optimization for personalized memory systems.
Figure 5: The left demonstrates CoMAM’s sensitivity to the credit assignment weight, and the right shows the credit assignment’s impact on each local agent’s performance on the 32K setting.
Table 5: Convergence step comparison of the joint and independent optimization.
Method Extract Profie Retrieval Total
32K
Independent 48 57 65 160
Joint 58 32 68 68
128K
Independent 210 280 300 790
Joint 300 350 320 350
1M
Independent 285 180 150 415
Joint 210 300 310 310
6 Conclusion
In this paper, we propose CoMAM, a collaborative reinforcement learning framework for jointly optimizing agents in personalized memory systems. It tackles the issue that independent agent optimization can not ensure global system performance without cross-agent collaboration. To enable joint optimization, we regularize the asynchronous executions of heterogeneous agents as a sequential MDP, which embeds inter-agent dependencies into state transitions and yields both local and global rewards. To align local improvement with global performance, we introduce an adaptive credit assignment mechanism that allocates global rewards to individual agents based on group-level ranking consistency between global and local rewards. Experiments on long-term personalized conversational benchmarks demonstrate that CoMAM outperforms leading memory baselines, validating the effectiveness of our collaborative RL framework. The limitation is discussed in Appendix E.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Algorithm
Here we present the algorithm of CoMAM’s training process in Algorithm 1.
Algorithm 1 Collaborative Reinforcement Learning Framework for Multi-Agent Memory Systems (CoMAM)
[…]
0: Optimized $\{\pi_{\theta}^{\text{cons}_{f}},\pi_{\theta}^{\text{cons}_{c}},\pi_{\theta}^{\text{ret}}\}$, response $p$
1:for $i=1$ to $G$ do
2:  $\mathcal{M}_{f}\leftarrow\pi_{\theta}^{\text{cons}_{f}}(\mathcal{H})$ $\triangleright$ Extract fine-grained memory
3:  $r_{\text{cons}_{f}}\leftarrow\alpha\cdot\frac{|\mathcal{M}_{f}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\alpha)\cdot\frac{|\mathcal{M}_{f}\cap\mathcal{E}_{q}|}{|\mathcal{M}_{f}\cup\mathcal{E}_{q}|}$
4:  $\mathcal{M}_{c}\leftarrow\pi_{\theta}^{\text{cons}_{c}}(\mathcal{M}_{f})$ $\triangleright$ Abstract coarse-grained memory
5:  $r_{\text{cons}_{c}}\leftarrow\mathcal{V}(\mathcal{M}_{c},\mathcal{H})$
6:  $\mathcal{M}\leftarrow\{\mathcal{M}_{f},\mathcal{M}_{c}\}$
7:  $(\mathcal{E}_{q}^{\prime},p)\leftarrow\pi_{\theta}^{\text{ret}}(\mathcal{M},q)$ $\triangleright$ Retrieve and generate response
8:  $r_{\text{ret}}\leftarrow\beta\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\beta)\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}^{\prime}\cup\mathcal{E}_{q}|}$
9:  $r_{\text{ans}}\leftarrow\mathbb{I}(p^{*}=p)$ $\triangleright$ Global reward (accuracy)
10:end for
11: Collect local reward $r_{n}^{(i)}$ for agent $n$, add to $\mathcal{R}_{\text{local},n}$
12: Collect global reward $r_{\text{ans}}^{(i)}$, add to $\mathcal{R}_{\text{global}}$
[…]
16:for $n=0$ to 22 do
17:  for $i=1$ to $G$ do
18:   $A_{n}^{(i)}\leftarrow\text{Normalize}(r_{\text{final},n}^{(i)})$ $\triangleright$ Compute advantage
19:   $\rho_{n}^{(i)}\leftarrow\frac{\pi_{\theta}^{n}(a_{n}^{(i)}\mid s_{n}^{(i)})}{\pi_{\theta_{\text{old}}}^{n}(a_{n}^{(i)}\mid s_{n}^{(i)})}$ $\triangleright$ Importance sampling ratio
20:  end for
21:  $\pi_{\theta}^{n}\leftarrow\text{Update}(\pi_{\theta}^{n},\mathcal{J}_{\text{GRPO}}(\pi_{\theta}^{n}))$ $\triangleright$ Optimize agent policy
22:end for
Appendix B Details of Experimental Settings
B.1 Dataset
In the PersonaMem benchmark, each interaction history consists of 10, 20, or 60 multi-turn sessions (corresponding to 32k, 128k, or 1M tokens) to form varying context lengths. Each session contains 15–30 conversation turns, designed to naturally reveal or update user preferences. In addition, each history is paired with a suite of subsequent queries across seven types, enabling comprehensive evaluation of personalized response capabilities. For PersonaMem’s three context length settings (32k, 128k, and 1M tokens), we partition their training and validation sets based on these queries. The statistical breakdown of training and validation queries is summarized in Table 6.
To process the training set, we first identify the specific session within the full conversational history that contains explicit evidence relevant to each query $q$, and we define this session as the query’s targeted conversational history $\mathcal{H}$. We then construct $(\mathcal{H},q)$ data pairs required for generating individual MDP trajectories as the core training data. The position of each target session within the complete interaction history is specified in the original benchmark. Ground-truth evidence $\mathcal{E}_{q}$ for each query $q$ is extracted using the Seed-1.6 model and validated through human evaluation. For evaluation, multiple distinct queries may share the same full conversational history that includes multiple sessions. This design enables further evaluation in complex inference scenarios where memory construction agents operate on a full conversational history with multiple sessions, while retrieval agents are activated for subsequent queries at varying frequencies.
B.2 Detailed Baselines
Here, we detailed our baseline methods, including methods without memory, prompt-based memory systems, and RL-optimized memory systems.
Methods without memory:
  • Base denotes the direct injection of raw conversational data as input to large language models (LLMs) for query answering tasks. Specifically, when the length of the input conversation exceeds the context window constraint, the input is truncated at the context window limit.
  • RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history via embedding-based similarity matching. The embedding model employed is all-MiniLM-L6-v2, and ultra-long input contexts are uniformly partitioned into fixed-length segments, each consisting of 2,048 tokens.
Prompt-based memory systems
  • A-Mem (Xu et al., 2025) instantiates a memory system through multiple agents with dedicated prompts, including the memory retrieval construction, precursor memory evolution, and retrieval agents;
  • CAM (Li et al., 2025) builds a hierarchical memory system via multiple agents, where dedicated agents undertake node construction (via ego-centric disentanglement and node replication) and information aggregation (for hierarchical abstraction). We implement it on the PersonaMem benchmark.
  • MemoryBank (Zhong et al., 2024) construct a human-like memory for LLMs that uses a forgetting curve to remember and forget information over time. Similarly, we implement it on the PersonaMem benchmark.
RL-based memory:
  • Mem1 (Zhou et al., 2026) proposes a reinforcement learning (RL) framework that empowers agents to conduct efficient multi-turn information retrieval and reasoning for query answering. The optimized agents prioritize direct memory retrieval from history, rather than memory construction.
  • Memory-R1 (Yan et al., 2025b) constructs a Memory Manager that learns optimal memory operations (ADD, UPDATE, DELETE) and an Answer Agent tasked with distilling relevant information from retrieved entries. Uniquely, the system is optimized via independent reinforcement learning (RL) in two stages: the Memory Manager is optimized first, followed by the Answer Agent.
B.3 More Implementation Details
For a fair comparison, all the baselines are implemented with the same LLM backbones as our CoMAM, including Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct. For the implementation of variant “+independent RL”, each agent is trained by fusing its task-specific rewards and global accuracy rewards, with the remaining agents frozen. For the “single policy” variant, we adopt a shared policy to serve as the policy for all three agents during training and inference. The prompts for different agents are presented below. The temperature of LLMs is set to $1$ for training and $0.8$ for evaluation. The detailed hyperparameter settings for CoMAM are presented in Table 7.
Table 6: The statistics of queries after the train-validation split across different dataset settings.
Dataset 32K 128K 1M
#Train 313 1,744 1,830
#Validate 100 200 200
#question types 7 7 7
Table 7: Hyperparameters of CoMAM on PersonaMem benchmark across different context lengths.
Dataset 32K 128K 1M
batch_size 128 128 128
group_size 8 8 8
max_model_length 32k 32k 32k
epochs 5 5 5
max_output_tokens 3k 3k 3k
Fine-grained memory construction – Extraction Agent You are the Memory Extractor. Your task is to analyze the upcoming Conversation History. Retain important conversations as much as possible in order to answer future questions about the conversation history, removing only the redundant text of each sentence. The conversation history is provided below: Conversation History:
   Historical Conversations
Output *only* the retained conversations from the conversation history you process.
Coarse-grained memory construction – Profile Agent You are the Profile Abstractor. You will receive fine-grained memory information from the Memory Extractor. Your task is to analyze the memory you receive and abstract it into higher-level knowledge for the user profile (such as scenarios, user behavior patterns, and preferences). Fine-grained Memory:
   Fine-grained Memory
Output only the abstracted user profile from the memory you receive.
Memory retrieval and query answer – Retrieval Agent You are the Retrieval agent. You have received fine-grained memory from the Memory Extractor and coarse-grained memory from the Profile Abstractor. Your task is to retrieve relevant memory to answer the following question accurately. Multi-grained Memory:
Fine-grained Memory from the Memory Extractor
   Coarse-grained memory from the Profile Abstractor
Question:
   User’s Subsequent Query
Options:
   The Four Options
Please output the retrieved relevant information in <information> </information> and choose the most appropriate option (a), (b), (c), or (d) in <final_answer></final_answer>.
Appendix C More Detailed Experimental Results
C.1 The Impact of Joint Optimization on Local Task Specialization
Joint optimization with adaptive credit assignment not only enhances global system performance but also promotes the specialization of each agent’s local tasks. As illustrated in Figure 6, we compare the local task reward curves of agents under three training strategies: optimizing with only local rewards (“w/ L”), optimizing with only global rewards (“w/ G”), and our joint optimization with adaptive credit assignment (“Ours”). Key observations are as follows:
  • Superior local reward convergence: For all three agents (Extraction, Profile, Retrieval), the local task rewards of “Ours” consistently outperform “w/ L” and “w/ G” in the stable phase. This indicates that adaptive credit assignment will not compromise local task capabilities; instead, it strengthens them by aligning local improvement with global goals.
  • Global-local Collaboration: The global reward curve of “Ours” converges to 0.70, significantly higher than “w/ L’ (0.65) and “w/ G” (0.64), proving that joint optimization realizes global performance enhancement by fostering inter-agent collaboration.
Adaptive credit assignment mechanism quantifies the correlation between each agent’s local performance and the global outcome, ensuring that agents receive rewards not only for completing local tasks well but also for contributing to the global goal. This guides agents to optimize their local strategies in a way that benefits the entire system, thus achieving both local specialization and global collaboration.
Figure 6: The curves for local task-specific reward $r_{n}$ and global accuracy reward $r_{3}$.
Appendix D Discussion
D.1 Difference with existing memory framework
Existing memory systems and our CoMAM differ fundamentally in optimization paradigm and reward design:
  • Optimization Paradigm: Independent Optimization vs. Joint Optimization
    Existing frameworks primarily adopt independent optimization. For example, MIRIX (Wang & Chen, 2025) designs a meta-agent and six specialized agents for memory management, optimizing each via tailored prompts; Memagent (Yu et al., 2026) focuses on a single retrieval agent, training it with query-answer accuracy as the reward; Memory-R1 (Yan et al., 2025b) splits the system into a Memory Manager and Answer Agent, training them in separate stages—first optimizing the Manager with the Agent frozen, then freezing the Manager to optimize the Agent. This approach ignores inter-agent collaboration: memory construction is not adjusted based on retrieval needs, and the Agent can only adapt to a fixed Manager’s output, leading to “local optimal but global suboptimal” outcomes.
    CoMAM proposes end-to-end joint optimization via MDP regularization. We model the asynchronous execution of Extraction, Profile, and Retrieval Agents as a sequential MDP trajectory, where each agent’s output directly serves as the next agent’s input. All three agents are updated simultaneously during training, with inter-agent dependencies embedded into state transitions, ensuring they evolve collaboratively toward the global goal of query-answer accuracy.
  • Reward Design: Single Global Reward vs. Adaptive Credit Assignment
    Existing frameworks that optimize memory agents with RL primarily rely solely on final query-answer accuracy as the reward for all agents. For example, Mem-$\alpha$ and Mem1 (Zhou et al., 2026) use this single reward to train construction and retrieval agents respectively. Memory-R1 (Yan et al., 2025b) and GAM (Yan et al., 2025a) also adopt this design, using only answer accuracy to optimize their agents separately. This ambiguous credit assignment prevents agents from distinguishing their individual contributions, leading to unstable convergence.
    CoMAM introduces adaptive integration of local and global rewards. We design task-specific local rewards (e.g., Extraction Agent’s information coverage, Profile Agent’s abstraction rationality, Retrieval Agent’s retrieval precision) and combine them with the global reward (answer accuracy). Using NDCG to quantify the ranking consistency between each agent’s local rewards and the global reward, we dynamically allocate global credit based on this consistency. This ensures agents are rewarded both for excelling at their local tasks and contributing to global performance, avoiding ambiguous optimization.
D.2 Difference with existing multi-agent system RL
  • Heterogeneous Policy vs. Homogeneous Policy: Some multi-agent reinforcement learning (RL) methods adopt a shared policy across all agents. For instance, CORY (Ma et al., 2024) proposes an RL framework featuring “Pioneer” and “Observer” agents that share a single policy model. These agents co-evolve through multi-agent collaboration and periodic role-swapping. In contrast, considering the heterogeneity and asynchrony of a memory system, CoMAM employs heterogeneous policies tailored to distinct agents.
  • Workflow: Sequential vs. Concurrent: Existing multi-agent system (MAS) methods often adopt a debate-style pipeline for reasoning tasks, such as math tasks. For instance, MAPoRL (Park et al., 2025) models agent interactions as concurrent debate, where multiple agents generate responses in parallel, focusing on persuasive collaboration for general reasoning. In contrast, CoMAM explicitly formulates agents’ executions as a sequential Markov Decision Process (MDP), embedding inter-agent dependencies into state transitions, ensuring that memory construction quality dynamically adapts to retrieval requirements and vice versa.
  • Collaboration Paradigm: Joint Optimization vs. Independent Alignment: Existing methods, such as Optimas (Wu et al., 2026), foster collaboration by aligning local rewards with global objectives via heuristic reward shaping for compound AI systems. However, it optimizes individual agents separately, leveraging locally-global aligned rewards rather than through end-to-end joint training. In contrast, CoMAM adopts a joint optimization paradigm: it regularizes the asynchronous execution of heterogeneous agents into a sequential MDP trajectory and updates all agents simultaneously.
  • Credit Assignment: Adaptive Assignment vs. Equal Allocation: Most existing systems use static credit assignment mechanisms without quantifying agents’ varying contributions: AT-GRPO (Zhao et al., 2026) distributes rewards equally across agents using fixed weights, ignoring task differences; CoMAS (Xue et al., 2026) adopts a zero-sum adversarial reward scheme that pits agents against each other, which is misaligned with the cooperative nature of memory systems; In contrast, CoMAM’s credit assignment is grounded in group-level ranking consistency (NDCG) between each agent’s local task rewards (e.g., coverage for extraction, rationality for profiling, precision for retrieval) and the global system reward (query-answer accuracy). This design quantifies each agent’s contribution to the global system. It dynamically adjusts weights according to agent performance, ensuring adaptive credit allocation that aligns local improvements with global performance.
Appendix E Limitation
While CoMAM enables effective joint optimization for multi-agent memory systems, it is worth noting that our current implementation focuses on verifying the core logic of collaborative RL, with certain extended capabilities reserved for future exploration to better adapt to more complex scenarios.
1. Limited Agent Task Scope:
The framework currently leverages a compact agent set (Extraction, Profile, Retrieval) to focus on verifying the efficacy of collaborative RL for personalized memory tasks. To prioritize the core verification goal, we reserve extended agent capabilities (such as memory modification for consistency maintenance and redundant memory deletion) for future work.
2. Static MDP Regularization for Training:
Current MDP regularization models agent execution as a static sequential trajectory, enabling joint training of heterogeneous and asynchronous agents. Compared with the existing independent optimization training paradigm, this design can better enhance the collaborative capabilities of agents beyond task-specific capabilities. While effective in boosting inter-agent collaboration, we can further consider the dynamic deployment of agents in real-world scenarios, such as dynamic invocation frequencies and real-time memory updates. Future work will integrate these real-world dynamics into the joint training process to improve the consistency between training and deployment.
3. Lack of Explicit local-local Alignment:
The adaptive credit assignment is specifically designed to align local agent improvements with global system goals, enabling agents’ collaboration towards the global goal. Since all agents’ local subtasks are oriented towards the consistent global system goal, the alignment between their local tasks is implicitly modeled. In the future, we will explore explicitly modeling this local-local alignment for agents’ local subtasks, which will further boost inter-agent collaboration.

### 27 Apr 2026 v2

@@
-Collaborative Multi-Agent Optimization for Personalized Memory System
+Joint Optimization of Multi-agent Memory System
 Abstract
-Memory systems are crucial to personalized LLMs by mitigating the context window limitation in capturing long-term user-LLM conversations.
-Typically, such systems leverage multiple agents to handle multi-granular memory construction and personalized memory retrieval tasks.
-To optimize the system, existing methods focus on specializing agents on their local tasks independently via prompt engineering or fine-tuning.
-However, they overlook cross-agent collaboration, where independent optimization on local agents hardly guarantees the global system performance.
-To address this issue, we propose a Collaborative Reinforcement Learning Framework for Multi-Agent Memory Systems (CoMAM), jointly optimizing local agents to facilitate collaboration.
-Specifically, we regularize agents’ execution as a sequential Markov decision process (MDP) to embed inter-agent dependencies into the state transition, yielding both local task rewards (e.g., information coverage for memory construction) and global rewards (i.e., query-answer accuracy).
-Then, we quantify each agent’s contribution via group-level ranking consistency between local and global rewards, treating them as adaptive weights to assign global credit and integrate local-global rewards.
-Each agent is optimized by these integrated rewards, aligning local improvements with the global performance.
-Experiments show CoMAM outperforms leading memory systems, validating the efficacy of our proposed collaborative reinforcement learning for joint optimization.
-Multi-agent, Memory System, Reinforcement Learning
+Memory systems are critical for LLMs, mitigating context window limitations and supporting long-horizon user–LLM interactions.
+Such systems typically comprise multiple agents responsible for memory construction and retrieval.
+Existing approaches often optimize each agent independently under a shared global objective (e.g., downstream QA accuracy), treating other agents as a static environment.
+However, this design has two key limitations: (1) independent optimization ignores inter-agent dependencies and lacks agents’ co-adaptation, and (2) relying solely on sparse global rewards provides limited guidance for optimizing specialized agents and causes ambiguous credit assignment.
+These may ultimately limit agents’ effective collaboration in the memory system.
+To address these limitations, we propose CoMAM, a joint optimization framework that promotes collaboration among agents via end-to-end reinforcement learning and an adaptive credit assignment mechanism.
+Specifically, we model the multi-agent pipeline as a Markov decision process (MDP) to expose inter-agent dependencies during end-to-end training.
+Agents are then jointly optimized using a combination of their local task reward and an adaptively weighted global reward, enabling agents to co-adapt while receiving targeted feedback for their respective roles.
+Experiments show that CoMAM consistently outperforms leading memory systems, validating the effectiveness of the joint optimization framework.
 1 Introduction
-Figure 1: Construction and retrieval agents, optimized on local tasks independently, yield lower global system performance than those under joint optimization.
-Memory systems (Chhikara et al., 2025; Zhong et al., 2024; Hu et al., 2025) have emerged as a pivotal frontier for personalized large language models (LLMs), which preserve users’ long-term conversational history beyond LLMs’ context window and selectively integrate it to tailor personalized responses (Chen et al., 2024a; Zhang et al., 2025).
-Typically, the memory system comprises multiple specialized agents: several construction agents summarize and store user history into multi-granular memories (e.g., long/short or coarse/fine), respectively (Li et al., 2025; Wang & Chen, 2025; Wang et al., 2025); and retrieval agents select and integrate relevant memories for subsequent queries via RAG (Lewis et al., 2020) or search-based reasoning (Yu et al., 2026; Jin et al., 2025).
-To enhance system effectiveness, existing methods focus on optimizing these agents for their respective tasks independently: Prompt engineering approaches (Li et al., 2025; Zhong et al., 2024) specialize each agent for its local task with well-crafted prompts, while reinforcement learning methods (Wang et al., 2025; Zhou et al., 2026; Yan et al., 2025b) train agents separately using rewards such as the query-answer accuracy.
-Despite local successes, they overlook cross-agent collaboration, treating the memory system as a “collection of local experts” rather than a collaborative team that evolves together for a global goal.
-Critically, independent optimization on local agents hardly guarantees global system performance (Akella, 2025; Wu et al., 2026).
-While individual agents may excel at their respective tasks, their independent strategies can conflict when combined in the global pipeline.
-To be more specific, if the fine-grained construction agent is optimized by treating retrieval agents as fixed environment components, it may store redundant or noisy content that impairs retrieval effectiveness.
-Conversely, if retrieval agents are trained without considering construction strategies, they may over-filter information and miss critical details during retrieval.
-As illustrated in Figure 1, even with independently well-trained agents (rewards converged), the combined system consistently underperforms that with joint optimization across different conversational history lengths.
-This motivates us to explore jointly optimizing multiple agents in the memory system to facilitate collaboration, which faces two challenges (see Figure 2):
-Figure 2: Illustration of challenges for joint optimization.
-  • Heterogeneity and Asynchrony: Multi-granular construction and retrieval agents operate with heterogeneous configurations and asynchronous executions across different tasks.
-Their discrete parameter space and inherent structural complexity complicate end-to-end optimization.
-  • Local-Global Alignment: Fostering collaboration requires aligning local improvement with the global performance.
-However, directly assigning equal shares of the global objective to individual agents introduces ambiguity in task execution, as agents’ contributions are divergent.
-To address these challenges, we introduce CoMAM, a Collaborative reinforcement learning framework for Multi-Agent Memory systems.
-Specifically, we formulate a standard memory system with three specialized agents for multi-granular memory construction and retrieval: an extraction agent to construct fine-grained memories (e.g., key events and facts), a profile agent to construct coarse-grained memories (e.g., users’ personalized preferences), and a retrieval agent to retrieve memories and answer queries.
-To optimize these heterogeneous agents jointly, we frame their asynchronous execution as a sequential Markov decision process (MDP), which embeds inter-agent dependencies into state transitions and yields both local task-specific rewards (e.g., information coverage for memory construction) and global system rewards (i.e., query answer accuracy).
-To achieve local-global alignment, CoMAM quantifies each agent’s system contribution by measuring group-level ranking consistency between its local and global rewards, then adaptively allocates global rewards to agents based on this consistency.
-Each agent is ultimately optimized with a combination of its task-specific reward and adaptively assigned global reward share, driving end-to-end collaboration toward the global objective.
-Extensive experiments on long-term personalized conversational tasks show CoMAM outperforms leading memory baselines, validating the efficacy of our collaborative reinforcement learning for jointly optimizing multi-agent memory systems.
+Memory systems Mem0 ; memorybank ; hu2025memory have emerged as a pivotal component for large language models (LLMs), enabling them to preserve long-horizon conversational history beyond the context window and selectively incorporate it for generation.
+Such systems are typically implemented as multi-agent pipelines, where construction agents summarize and store user history into multi-granular memories (e.g., long/short or coarse/fine) li2025cam ; mirix ; memory-alpha , and retrieval agents select and integrate relevant memories for downstream queries via RAG RAG or search-based reasoning memagent ; search-r1 .
+These agents are often instantiated with heterogeneous configurations to handle tasks of varying complexity, enabling flexible modular design and specialization.
+To enhance the performance of memory systems, existing approaches mem-r1 ; mem1 ; GAM ; memory-alpha often optimize each agent independently via reinforcement learning under a shared global objective (e.g., , QA accuracy on downstream tasks).
+They typically update one agent at a time, treating the others as fixed components of the environment.
+However, such independent optimization has two limitations and may lead to sub-optimal division of labor and collaboration local_optimalty ; Optimas .
+First, ignorance of dependency among agents: each agent is optimized while treating others as a static environment, preventing the adaptation to other agents’ evolving policies.
+Second, ambiguous credit assignment: the shared global objective provides insufficient signal to attribute performance improvements to individual agents, which may hinder effective specialization.
+As illustrated in Figure 1, although agents can be well trained independently (rewards converged in the left figure), combining them directly leads to inferior system performance compared to joint optimization across different context lengths (lower QA accuracy in the middle figure).
+This gap highlights that independent optimization with the same global system objective is insufficient to achieve optimal system performance, motivating us to jointly optimize these heterogeneous agents in the memory system to facilitate collaboration.
+Figure 1: (Left) Training curves of coarse/fine memory construction and retrieval agents, which are optimized independently.
+(Middle) Global system performance under independent and joint optimization across different context lengths.
+(Right) Conceptual comparison between existing methods that optimize agents independently with shared global rewards and our joint optimization with adaptive credit assignment.
+To address these limitations, we propose CoMAM, a joint optimization framework that promotes collaboration among agents via end-to-end reinforcement learning, together with an adaptive credit assignment mechanism.
+As illustrated in the right panel of Figure 1, our approach differs from independent optimization by enabling simultaneous updates of heterogeneous agents with adaptively assigned rewards.
+Specifically, to formulate a standard memory system for coarse/fine memory construction and memory retrieval, we design an extraction agent to construct fine-grained memories (e.g., key events and facts), a profile agent to summarize coarse-grained memories (e.g., users’ stable preferences), and a retrieval agent to retrieve memories and answer queries.
+To model the inter-agent dependencies for co-adaption, we frame the heterogeneous multi-agent pipeline as a sequential Markov decision process (MDP), enabling agents to coordinate through their intermediate outputs.
+To alleviate ambiguous credit assignment during joint optimization, CoMAM assigns each agent a combination of its local task reward and an adaptively weighted global reward, providing more targeted feedback for different roles, driving collaboration for the global objective while maintaining task specialization.
+Extensive experiments on long-horizon conversational tasks show CoMAM outperforms leading memory baselines, validating the efficacy of our joint optimization framework and adaptive credit assignment for multi-agent memory systems.
 2 Related work
-Memory System for Personalized LLMs
+Memory Systems for LLMs.
 […]
-Optimization of Multi-agent System
-Multi-agent Systems (MAS) (Sun et al., 2024; Chen et al., 2024b) leverage multiple specialized agents with heterogeneous configurations and asynchronous execution to tackle complex tasks beyond the capacity of a single agent.
-To boost MAS performance, existing methods focus on independent optimization, primarily via prompt engineering (Patil et al., 2024; Wu et al., 2024) or fine-tuning of agents on specific tasks (Subramaniam et al., 2025; Yan et al., 2025b).
-Yet such methods yield only marginal gains, as they overlook the value of joint optimization for inter-agent collaboration.
-Consequently, recent research has shifted to reinforcement learning (RL)-based collaborative training for end-to-end system optimization, such as MAPoRL (Park et al., 2025) and MARFT (Liao et al., 2025).
-Credit Assignment
-Credit assignment(Zhou et al., 2020; Bo et al., 2024a; Tseng et al., 2022) concerns properly attributing a team’s global outcome to the specific agents in Multi-Agent Reinforcement Learning (MARL).
-This is particularly difficult in quantifying agents’ contributions to the outcome, which is often obscured by the actions of other agents and environmental noise.
-Naive credit assignment approaches evenly distribute global team rewards across agents (Zhao et al., 2026; Hong et al., 2025).
-Though simple, this leads to significant credit assignment ambiguity: agents cannot distinguish their specific impact, resulting in high variance in policy updates and unstable convergence.
-3 Preliminary and Challenges
+Optimization of Multi-agent Systems.
+Multi-agent systems (MAS) masrl_syrvey ; self_play ; Optimas leverage multiple specialized agents with heterogeneous configurations to tackle complex tasks beyond the capacity of a single agent.
+Existing approaches typically optimize agents independently, either through prompt engineering Gorilla ; avatar or task-specific fine-tuning independent_mas ; mem-r1 .
+Such independent optimization treats other agents as part of the environment and does not explicitly account for their inner dependence during training.
+Consequently, recent research has shifted to RL-based joint training for end-to-end system optimization, such as MAPoRL MAPORL and MARFT MARFT .
+Credit Assignment.
+Credit assignment credit_assign3 ; credit_assign5 ; credit_assign4 concerns properly attributing a team’s global outcome to the specific agents in Multi-Agent Reinforcement Learning.
+This is challenging due to the interdependence among agents and environmental noise, which obscures each agent’s contribution.
+A common strategy is to distribute global rewards uniformly across agents naive_assign ; hong2025multi .
+While simple, this leads to significant credit assignment ambiguity: agents cannot distinguish their specific contributions COMA ; credit_assignment3 , resulting in high variance in policy updates and unstable convergence.
+3 Preliminary
 3.1 Task Formulation for Multi-agent Memory System
-We define the user’s long-term conversational history as $\mathcal{H}=\{h_{1},\dots,h_{N}\}$, which consists of $N$-turn interactions that exceed the LLM’s context window.
-To facilitate personalization, the Memory System transforms the history $\mathcal{H}$ into multi-granular memories $\mathcal{M}$ that can be retrieved to answer users’ subsequent queries $q$, realized by diverse agents on construction and retrieval tasks (Hu et al., 2025):
-  • Memory Construction: The agents (Li et al., 2025) under policies $\{\pi_{\theta}^{\text{cons}_{g}}\}$, distill the raw conversational history $\mathcal{H}$ at different granularities $g$, respectively, finally constructing a multi-granular memory set $M$:
-    $\mathcal{M}_{g}=\pi_{\theta}^{\text{cons}_{g}}(\mathcal{H}),$ (1)
-    $\mathcal{M}=\{\mathcal{M}_{g}\}.$ (2)
-  • Memory Retrieval: Given a subsequent query $q$, retrieval agents (Zhou et al., 2026) under policies $\{\pi_{\theta}^{\text{ret}}\}$ identify the $K$ most relevant memories from $\mathcal{M}$.
+We define the long-horizon user-LLM conversational history as $\mathcal{H}=\{h_{1},\dots,h_{N}\}$, which consists of $N$-turn interactions that exceed the LLM’s context window.
+The Memory System transforms the history $\mathcal{H}$ into multi-granular memories $\mathcal{M}$ that can be retrieved to answer users’ subsequent queries $q$, realized by heterogeneous agents on construction and retrieval tasks hu2025memory :
+  • Memory Construction: The agents li2025cam under policies $\{\pi_{\theta}^{\text{cons}_{g}}\}$, distill the raw conversational history $\mathcal{H}$ at different granularities $g$, finally constructing a multi-granular memory set $M$:
+    $\mathcal{M}_{g}=\pi_{\theta}^{\text{cons}_{g}}(\mathcal{H}),\quad\mathcal{M}=\{\mathcal{M}_{g}\}.$ (1)
+  • Memory Retrieval: Given a subsequent query $q$ of downstream task, retrieval agents mem1 under policies $\{\pi_{\theta}^{\text{ret}}\}$ identify the $K$ most relevant memories from $\mathcal{M}$.
 Then, the agent with policy $\pi_{\theta}^{\text{LLM}}$ generates a response $p$ for the query $q$ using the retrieved $K$ memories.
 Formally:
-    $\{\bm{m}_{k}\}_{k=1}^{K}=\pi_{\theta}^{\text{ret}}(q,\mathcal{M},K),$ (3)
-    $p=\pi_{\theta}^{\text{LLM}}\left(q,\{\bm{m}_{k}\}_{k=1}^{K}\right).$ (4)
-    For efficiency, $\pi_{\theta}^{\text{ret}}$ and $\pi_{\theta}^{\text{LLM}}$ can be unified into a single policy that performs both memory retrieval and personalized response generation for query $q$.
-The global goal of the memory system is to align the generated response $p$ with the user’s personalized preferences and experiences.
-Formally, given a dataset $\mathcal{D}$ with input pairs $(\mathcal{H},q)$ and a performance evaluation function $\mathcal{F}$, the optimization objective is to find the optimal policy set $\Pi^{*}$ that maximizes the expected global performance:
-$\Pi^{*}=\arg\max_{\Pi}\mathbb{E}_{(\mathcal{H},q)\sim\mathcal{D}}\left[\mathcal{F}\left(p^{*},\Pi(\mathcal{H},q)\right)\right].$ (5)
-Here, $\Pi=\{\pi_{\theta}^{\text{cons}_{g}},\pi_{\theta}^{\text{ret}}\}$ denotes the policy set, and $p^{*}$ denotes the ground-truth response.
+    $\{\bm{m}_{k}\}_{k=1}^{K}=\pi_{\theta}^{\text{ret}}(q,\mathcal{M},K),\quad p=\pi_{\theta}^{\text{LLM}}\left(q,\{\bm{m}_{k}\}_{k=1}^{K}\right).$ (2)
+    For efficiency, $\pi_{\theta}^{\text{ret}}$ and $\pi_{\theta}^{\text{LLM}}$ can be unified into a single policy that performs both memory retrieval and response generation for query $q$.
+The goal of the memory system is to optimize downstream task performance.
+Given a dataset $\mathcal{D}$ of input pairs $(\mathcal{H},q)$ and an evaluation function $\mathcal{F}$, we aim to learn an optimal policy set $\Pi^{*}$ that maximizes the expected performance:
+$\Pi^{*}=\arg\max_{\Pi}\mathbb{E}_{(\mathcal{H},q)\sim\mathcal{D}}\left[\mathcal{F}\big(p^{*},\Pi(\mathcal{H},q)\big)\right].$ (3)
+Here, $\Pi=\{\pi_{\theta}^{\text{cons}_{g}},\pi_{\theta}^{\text{ret}}\}$ denotes the policy set, and $p^{*}$ is the ground-truth response.
 3.2 Reinforcement Learning for LLMs
-To optimize LLMs’ policy $\pi_{\theta}$, Reinforcement Learning (RL) (Ouyang et al., 2022; Zheng et al., 2023) is typically used to maximize the expected reward while maintaining stability relative to a reference model $\pi_{\text{ref}}$.
-In this work, we primarily build upon Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an efficient RL algorithm that eliminates the need for a separate value function (critic) through group-based relative rewards.
+To optimize LLMs’ policy $\pi_{\theta}$, Reinforcement Learning (RL) RLHF ; PPO is typically used to maximize the expected reward while maintaining stability relative to a reference model $\pi_{\text{ref}}$.
+In this work, we primarily build upon Group Relative Policy Optimization (GRPO) GRPO , an efficient RL algorithm that eliminates the need for a separate value function (critic) through group-based relative rewards.
 Given the input $x$, GRPO samples a group of outputs $\left\{o_{1},\ldots,o_{G}\right\}$ from the current policy $\pi_{\theta}$ and optimizes the objective function:
-$\mathcal{J}_{GRPO}\left(\pi_{\theta}\right)=$ (6)
-$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left\{\min\left[\rho_{i,t}\hat{A}_{i,t},\hat{\rho}_{i,t}\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}\right\},$
+$\mathcal{J}_{GRPO}\left(\pi_{\theta}\right)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\sum_{t=1}^{\left|o_{i}\right|}\left\{\min\left[\rho_{i,t}\hat{A}_{i,t},\hat{\rho}_{i,t}\hat{A}_{i,t}\right]-\beta\mathbb{D}_{KL}\right\},$ (4)
 where $\rho_{i,t}=\frac{\pi_{\theta}\left(o_{i,t}\mid x,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid x,o_{i,<t}\right)}$ is the importance sampling ratio, $\hat{\rho}_{i,t}=\operatorname{clip}\left(\rho_{i,t};1-\epsilon,1+\epsilon\right)$ is the clipped ratio.
@@
 Formally, we have:
-$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)}.$ (7)
-The KL divergence is defined as:
-$\mathbb{D}_{KL}=\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-1.$ (8)
+$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)},\text{ and }\quad\mathbb{D}_{KL}=\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|x,o_{i,<t})}{\pi_{\theta}(o_{i,t}|x,o_{i,<t})}-1.$ (5)
 4 Method
-Figure 3: The overview of our proposed framework, CoMAM, which regularizes agents’ execution as MDP trajectories for joint RL optimization and fosters collaboration via adaptive credit assignment to achieve local-global alignment.
-In this section, we present our proposed CoMAM, a Collaborative Reinforcement Learning Framework designed to address the key challenges of jointly optimizing the multi-agent memory systems, as shown in Figure 3.
+Figure 2: The overview of our joint optimization framework for the multi-agent memory system, CoMAM, which promotes collaboration among agents via end-to-end reinforcement learning and an adaptive credit assignment mechanism.
+In this section, a joint optimization framework that enables agents to co-adapt via end-to-end reinforcement learning and an adaptive credit assignment mechanism, as shown in Figure 2.
 We begin by defining the task-specific agents alongside the local and global rewards in Sec 4.1.
-Then, we elaborate on regularizing the execution trajectory as a Markov decision process (MDP) in Sec 4.2, which enables joint optimization despite agent heterogeneity and asynchrony.
-Finally, we present the adaptive credit assignment mechanism in Sec 4.3, which quantifies individual agent contributions and aligns local improvements with global performance.
+Then, we frame the heterogeneous multi-agent pipeline as a Markov decision process (MDP) in Sec 4.2, which models the inter-agent dependencies for joint training.
+Finally, we present the adaptive credit assignment mechanism in Sec 4.3, which quantifies individual agent contributions and provides more targeted feedback for different agents.
 4.1 Task-Specific Agents Definition
-Building on existing multi-agent memory paradigms (Hu et al., 2025; Li et al., 2025), we define a standard memory system consisting of three specialized agents: Extraction Agent, Profile Agent, and Retrieval Agent.
+Building on existing multi-agent memory paradigms hu2025memory ; li2025cam , we define a standard memory system consisting of three specialized agents: Extraction Agent, Profile Agent, and Retrieval Agent.
 Extraction Agent extracts fine-grained memory $\mathcal{M}_{f}$ with policy $\pi_{\theta}^{\text{cons}_{f}}$ from the raw conversational history $\mathcal{H}$.
@@
 $r_{\text{cons}_{f}}=\alpha\cdot\frac{|{\mathcal{M}_{f}}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\alpha)\cdot\frac{|{\mathcal{M}_{f}}\cap\mathcal{E}_{q}|}{|{\mathcal{M}_{f}}\cup\mathcal{E}_{q}|},$ (6)
-where $r_{\text{cons}_{f}}$ balances the coverage of relevant information and the filtering of irrelevant content. To prioritize coverage, $\alpha$ is set to $0.8$.
+where $r_{\text{cons}_{f}}$ balances the coverage of relevant information and the filtering of irrelevant content.
+To prioritize fact coverage during the memory construction stage, $\alpha$ is set to $0.8$.
 Profile Agent abstracts coarse-grained memory $\mathcal{M}_{c}$ (e.g., user’s preferences or behavioral patterns) from $\mathcal{M}_{f}$, guided by policy $\pi_{\theta}^{\text{cons}_{c}}$.
@@
 where $r_{\text{cons}_{c}}$ evaluates the quality of $\mathcal{M}_{c}$ relative to the conversation history $\mathcal{H}$ according to predefined rubrics (e.g., scenario, preference, and user profile).
+More details are presented in Appendix B.3.
 $\mathcal{V}$ denotes a frozen pre-trained LLM for reward generation.
 Finally, the comprehensive memory set is defined as $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$.
-Retrieval Agent retrieves top-$k$ relevant memories $\mathcal{E}_{q}^{\prime}=\{\bm{m}_{k}\}_{k=1}^{K}$ from memory set $\mathcal{M}$ and generate personalized response $p$ for user’s subsequent query $q$, guided by policy $\pi_{\theta}^{\text{ret}}$.
+Retrieval Agent retrieves top-$k$ relevant memories $\mathcal{E}_{q}^{\prime}=\{\bm{m}_{k}\}_{k=1}^{K}$ from memory set $\mathcal{M}$ and generate response $p$ for user’s subsequent query $q$, guided by policy $\pi_{\theta}^{\text{ret}}$.
 To evaluate the retrieval performance, we define a rule-based reward function that enables precise matching between the recalled $\mathcal{E}_{q}^{\prime}$ and ground-truth evidence $\mathcal{E}_{q}$:
 $r_{\text{ret}}=\beta\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\beta)\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}^{\prime}\cup\mathcal{E}_{q}|},$ (8)
-where $r_{\text{ret}}$ balances the coverage and precision of retrieval. To prioritize retrieval precision while minimizing noise, $\beta$ is set to $0.2$.
+where $r_{\text{ret}}$ balances the coverage and precision of retrieval. To prioritize retrieval precision while minimizing noise for answer generation, $\beta$ is set to $0.2$.
 To evaluate the memory system’s global performance, we define the rule-based reward for $\pi_{\theta}^{\text{ret}}$ to evaluate its answering accuracy for query $q$.
-Formally, let $p^{*}$ denote the ground-truth response for the personalized query $q$, we have:
+Formally, let $p^{*}$ denote the ground-truth response for the query $q$, we have:
 $r_{\text{ans}}=\mathbb{I}\left(p^{*}=p\right)$ (9)
@@
 Notably, the number and tasks of these agents can be further expanded to meet diverse application requirements in a multi-agent memory system.
-4.2 Execution Trajectory Regularization
-Usually, construction agents run periodically to manage historical conversations, while retrieval agents are activated for incoming queries.
-To enable joint training under such heterogeneous and asynchronous settings, we model the entire construction and retrieval pipeline of the multi-agent memory system as a sequential Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},T,R)$, where each step is executed by a heterogeneous agent.
-Formally, let $\pi_{\theta}^{0}=\pi_{\theta}^{\text{cons}_{f}}$, $\pi_{\theta}^{1}=\pi_{\theta}^{\text{cons}_{c}}$$\pi_{\theta}^{2}=\pi_{\theta}^{\text{ret}}$ denote the policies at step $n=1,2,3$, respectively.
-Then, for each MDP trajectory $\tau$, we have:
-  • $\mathcal{S}$: State space, capturing the system state at each step in the MDP.
-Specifically, $s_{0}=\mathcal{H}$ denotes the Extraction Agent’s initial state of raw conversational history $\mathcal{H}$.
-$s_{1}=\mathcal{M}_{f}$ denotes the Profile Agent’s state, where $\mathcal{M}_{f}$ is the fine-grained memory extracted by the prior Extraction Agent.
-$s_{2}=(\mathcal{M},q)$ denotes the Retrieval Agent’s state, with $\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{c}\}$ as the multi-granular memory set and $q$ as the incoming user query.
-$s_{3}=(\mathcal{E}_{q}^{\prime},p)$ denotes the system’s final state, comprising retrieved information $\mathcal{E}_{q}$ and generated response $p$.
-  • $\mathcal{A}$: Action space, defining executable actions of each heterogeneous agent in the MDP.
-Specifically, $a_{0}$ denotes extracting fine-grained memory by the Extraction Agent.
-$a_{1}$ denotes abstracting coarse-grained memory by the Profile Agent.
-$a_{2}$ denotes the Retrieval Agent’s action, which involves retrieving top-$k$ relevant memories and generating a personalized response.
-  • $T$: State transition function, where $s_{n+1}=T(s_{n},a_{n})$ and $a_{n}=\pi_{\theta}^{n}(s_{n})$.
-Specifically, the Extraction Agent transitions to the Profile Agent via $s_{1}=T(s_{0},a_{0})$ and $a_{0}=\pi_{\theta}^{0}(\mathcal{H})$; the Profile Agent then transitions to the Retrieval Agent through $s_{2}=T(s_{1},a_{1})$ and $a_{1}=\pi_{\theta}^{1}(\mathcal{M}_{f})$; finally, the Retrieval Agent produces the system’s final state $s_{3}=T(s_{2},a_{2})$ where $a_{2}=\pi_{\theta}^{2}(\mathcal{M},q)$.
-  • $R$: Reward function, quantifying the rewards of agent actions in the MDP.
-As detailed in Sec 4.1, the local task rewards for each step are $r_{0}=r_{\text{cons}_{f}}$ for the Extraction Agent, $r_{1}=r_{\text{cons}_{c}}$ for the Profile Agent, and $r_{2}=r_{\text{ret}}$ for the Retrieval Agent.
-And, the global system reward is defined as $r_{3}=r_{\text{ans}}$.
+4.2 Inter-agent Dependency Modeling
+In a typical memory system, construction agents first convert conversational histories into structured memories, which are then consumed by retrieval agents to answer queries.
+This naturally forms a sequential pipeline for heterogeneous agents to coordinate through their intermediate outputs RL_MDP .
+To capture such inter-agent dependencies during co-adaptation, we model the pipeline as a sequential Markov decision process (MDP), enabling joint optimization across heterogeneous agents.
+This formulation serves as a training-time abstraction: it does not assume synchronous deployment, but provides a structured way to expose how upstream decisions influence downstream behavior.
+Formally, the MDP is defined as $(\mathcal{S},\mathcal{A},T,R)$, where each step corresponds to a heterogeneous agent executing its policy $\{\pi_{\theta}^{n}\}_{n=0}^{2}$ (extraction, profile, retrieval).
+Given an input $(\mathcal{H},q)$, the system evolves as a trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},s_{2},a_{2})$:
+  • $\mathcal{S}$: The state represents intermediate outputs of each agent in the MDP, with $s_{0}=\mathcal{H}$, $s_{1}=\mathcal{M}_{f}$, $s_{2}=(\mathcal{M},q)$, and $s_{3}=(\mathcal{E}_{q}^{\prime},p)$.
+  • $\mathcal{A}$: Each action $a_{n}$ corresponds to the execution of agent $\pi_{\theta}^{n}$, including fine-grained memory extraction, coarse-grained abstraction, and retrieval with response generation.
+  • $T$: The transition function $s_{n+1}=T(s_{n},a_{n})$ propagates intermediate outputs, where each agent’s output becomes part of the next agent’s input, explicitly exposing inter-agent dependencies.
+  • $R$: Each step is associated with a local task reward $r_{n}$, together with a shared global reward $r_{3}$ reflecting downstream performance.
 Within this formulation, CoMAM embeds inter-agent dependencies into state transitions along the optimization trajectory.
-The multi-agent memory system operates sequentially to generate local and global rewards, enabling end-to-end RL optimization with simultaneous updates of heterogeneous policies.
-The task-specific capabilities and inter-agent dependencies learned in this MDP enable agents to perform well in more complex inference environments, where divergent operating frequencies demand enhanced collaboration.
+The multi-agent memory system operates sequentially to generate local and global rewards, enabling end-to-end RL for heterogeneous policies’ co-adaptation.
 4.3 Adaptive Credit Assignment Mechanism
-To align local improvements with global performance during joint optimization, we propose an adaptive credit assignment mechanism to allocate global system rewards to local agents.
-Instead of assigning global rewards equally to each agent (Hong et al., 2025; Zhao et al., 2026), CoMAM quantifies agents’ varying contributions to the system and treats them as adaptive weights for allocation.
-Formally, for each input $(\mathcal{H},q)$, we sample a group of MDP trajectories $\{\tau_{i}\}_{i=1}^{G}$, collecting the local rewards for each agent $n$, $\mathcal{R}_{\text{local},n}=\{r_{n}^{(i)}\}_{i=1}^{G}$, and the corresponding global rewards, $\mathcal{R}_{\text{global}}=\{r_{3}^{(i)}\}_{i=1}^{G}$, where $n\in\{0,1,2\}$.
-To quantify agent $n$’s contribution to the memory system, we compute the ranking consistency score between $\mathcal{R}_{\text{local},n}$ and $\mathcal{R}_{\text{global}}$ using the Normalized Discounted Cumulative Gain (NDCG) metric.
-Specifically, the NDCG-based consistency score for agent $n$ is defined as:
+While modeling inter-agent dependencies during joint training enables co-adaptation, relying solely on a shared global reward provides limited and ambiguous guidance for optimizing specialized agents hong2025multi ; naive_assign .
+In particular, when agents cannot distinguish their individual impact on the overall outcome, they receive indistinguishable learning signals, which prevent the development of effective role-specific policies and consequently hinder agents’ collaboration in the memory system.
+To address this limitation, we introduce an adaptive credit assignment mechanism that distributes the global reward among agents in proportion to their estimated contribution, providing more informative training signals for each specialized agent.
+Formally, for each input $(\mathcal{H},q)$, we sample a group of MDP trajectories $\{\tau_{i}\}_{i=1}^{G}$, and collect local rewards for each agent $n$, $\mathcal{R}_{\text{local},n}=\{r_{n}^{(i)}\}_{i=1}^{G}$, along with the corresponding global rewards, $\mathcal{R}_{\text{global}}=\{r_{3}^{(i)}\}_{i=1}^{G}$, where $n\in\{0,1,2\}$.
+To estimate each agent’s contribution to the global system, we measure the alignment between its local and global rewards using a ranking-based consistency metric, instantiated as Normalized Discounted Cumulative Gain (NDCG) NDCG_1 :
 $v_{n}=\text{NDCG}(\sigma(\mathcal{R}_{\text{local},n}),\sigma(\mathcal{R}_{\text{global}})),$ (10)
-where the global reward ranking $\sigma(\mathcal{R}_{\text{global}})$ serves as the ideal ranking for the local reward ranking $\sigma(\mathcal{R}_{\text{local},n})$.
-We then normalize $v_{n}$ to obtain adaptive credit weights $w_{n}$ that satisfy $\sum_{n}w_{n}=1$:
-$w_{n}=\frac{\exp(v_{n})}{\sum_{n^{\prime}}\exp(v_{n^{\prime}})}.$ (14)
-The final optimization signal for agent $n$ is the combination of its local rewards and allocated global credit, weighted by the adaptive weight $w_{n}$:
+where $\sigma(\cdot)$ denotes the ranking operator.
+Intuitively, this metric captures whether trajectories with higher global rewards are also associated with higher local rewards for a given agent.
+Agents whose local reward rankings are more consistent with global outcome rankings receive higher credit, as their behaviors are more aligned with improving overall performance.
+Moreover, compared to value-based metrics, ranking-based measures are more robust under sparse and noisy rewards, as they rely on relative ordering rather than absolute magnitudes ranking_reward .
+The resulting scores are normalized to obtain adaptive weights $w_{n}$ with $\sum_{n}w_{n}=1$: $w_{n}=\frac{\exp(v_{n})}{\sum_{n^{\prime}}\exp(v_{n^{\prime}})}$.
+Each agent’s final optimization signal combines its local reward with its adaptively weighted share of the global reward:
 $r_{\text{final},n}^{(i)}=r_{n}^{(i)}+w_{n}\cdot r_{3}^{(i)}.$ (11)
-After the adaptive credit assignment, we apply the GRPO algorithm (as defined in Sec 3.2) on each agent based on the integrated rewards $r_{\text{final}}$.
-Formally, the objective function for the agent $n$ is defined as:
-$\mathcal{J}_{GRPO}\left(\pi_{\theta}^{n}\right)=$ (16)
-$\frac{1}{G}\sum_{i=1}^{G}\min\left[\rho_{n}^{(i)}A_{n}^{(i)},\;\text{clip}\left(\rho_{n}^{(i)},\;1\pm\epsilon\right)A_{n}^{(i)}\right],$
-where $\rho_{n}^{(i)}=\frac{\pi_{\theta}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}{\pi_{\theta_{\text{old}}}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}$ is the importance sampling ratio, the advantage $A_{n}^{(i)}$ for each agent’s action $a_{n}^{(i)}$ is computed by normalizing its reward $r_{\text{final},n}^{(i)}$.
-By optimizing each agent with both local task-specific rewards and allocated global system rewards, we can align local improvements with global performance and foster collaboration.
+We then optimize each agent simultaneously using the GRPO algorithm (as introduced in Sec.
+3.2) on these integrated rewards: $\mathcal{J}_{\text{GRPO}}\left(\pi_{\theta}^{n}\right)=\frac{1}{G}\sum_{i=1}^{G}\min\Big[\rho_{n}^{(i)}A_{n}^{(i)},\;\text{clip}(\rho_{n}^{(i)},1\pm\epsilon)A_{n}^{(i)}\Big]$, where $\rho_{n}^{(i)}$ is the importance sampling ratio and $A_{n}^{(i)}$ the normalized advantage computed from $r_{\text{final},n}^{(i)}$.
+By leveraging both local task-specific rewards and adaptively weighted global rewards, this mechanism fosters collaboration while preserving agent specialization.
 5 Experiments
 In this section, we conduct extensive experiments to validate the effectiveness of our proposed CoMAM by addressing the following questions:
-  • RQ1: How does CoMAM perform against leading baselines of memory systems for personalization?
+  • RQ1: How does CoMAM perform against leading memory systems?
   • RQ2: What is the contribution of each agent within the memory system?
-  • RQ3: What is the impact of regularizing agents’ execution as an MDP for jointly optimizing the memory system?
-  • RQ4: What is the contribution of adaptive credit assignment for facilitating collaboration in the memory system?
+  • RQ3: What is the impact of inter-agent dependency modeling for joint optimization?
+  • RQ4: What is the contribution of adaptive credit assignment for the specialization agents?
   • RQ5: What is the computational cost of the CoMAM?
-Table 1: Query-Answering Accuracy on the PersonaMem Benchmark across 32K, 128K, and 1M context length settings. Bold indicates the best performance, and underlining denotes the second best.
+Table 1: Query-Answering Accuracy on the PersonaMem Benchmark across 32K, 128K, and 1M context lengths. Bold indicates the best performance, and underlining denotes the second best.
 Method Qwen Llama
@@
 Base 0.41 0.39 0.38 0.35 0.31 0.32
-RAG (Lewis et al., 2020) 0.48 0.45 0.41 0.43 0.39 0.36
+RAG RAG 0.48 0.45 0.41 0.43 0.39 0.36
 Prompt-based memory
-CAM (Li et al., 2025) 0.53 0.50 0.45 0.48 0.45 0.43
-MemoryBank (Zhong et al., 2024) 0.51 0.47 0.44 0.45 0.42 0.39
-A-Mem (Xu et al., 2025) 0.50 0.44 0.45 0.46 0.37 0.36
+CAM li2025cam 0.53 0.50 0.45 0.48 0.45 0.43
+MemoryBank memorybank 0.51 0.47 0.44 0.45 0.42 0.39
+A-Mem xu2025amem 0.50 0.44 0.45 0.46 0.37 0.36
 RL-based memory
-Mem1 (Zhou et al., 2026) 0.59 0.57 0.58 0.56 0.58 0.61
-Memory-R1 (Yan et al., 2025b) 0.58 0.60 0.60 0.57 0.61 0.60
+Mem1 mem1 0.59 0.57 0.58 0.56 0.58 0.61
+Memory-R1 mem-r1 0.58 0.60 0.60 0.57 0.61 0.60
 Ours 0.64 0.70 0.66 0.62 0.68 0.69
 Impro +8.5% +16.7% +10% +8.8% +11.5% +13.1%
+Table 2: Results on LongMemEval using Qwen2.5-7B.
+Method $F_{1}\uparrow$ $B_{1}\uparrow$ $J_{1}\uparrow$
+RAG 18.27 14.57 22.20
+SeCom 19.03 16.57 10.00
+A-Mem 41.55 36.58 54.80
+Mem0 38.44 34.53 46.80
+GAM 51.75 48.96 52.00
+Ours 62.17 51.88 57.60
 5.1 Experimental Settings
 Dataset
-We conduct experiments on the PersonaMem benchmark (Jiang et al., 2025), a large-scale dataset including over 180 long-term user-LLM interaction histories and about 6,000 multiple-choice questions as subsequent queries.
-The benchmark is split into three experimental settings with varying history lengths (i.e., 32K, 128K, and 1M tokens), each paired with subsequent queries spanning seven distinct types: (1) Recall user-shared facts, (2) Suggest new ideas, (3) Acknowledge latest user preferences, (4) Track full preference evolution, (5) Revisit reasons for preference updates, (6) Provide preference-aligned recommendations, and (7) Generalize to new scenarios.
-Compared to existing benchmarks (Maharana et al., 2024; Wu et al., 2025), PersonaMem’s strengths in long context history and personalized query type enable rigorous assessment of memory systems in personalized conversational scenarios.
+We conduct experiments on the PersonaMem personamem and LongMemEval longmemeval benchmarks.
+PersonaMem provides long-horizon conversational histories with varying context lengths (i.e., 32K, 128K, and 1M tokens), each paired with seven downstream query types.
+LongMemEval focuses on multi-session conversational settings, requiring tracking information and performing temporal reasoning across sessions.
+They offer a more realistic evaluation of long-horizon memory construction and retrieval.
 Dataset processing details are provided in Appendix B.1.
@@
 We conduct a comprehensive evaluation of CoMAM against multiple leading baselines, including methods without a memory system “No memory”, methods of building memory systems via dedicated prompts “Prompt-based memory”, and methods of optimizing the memory systems with RL “RL-based memory”.
-Detailed introductions and implementations of baselines are provided in Appendix B.
-  • No memory: “Base” represents directly injecting raw conversations as LLMs’ input for query answering; RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history;
-  • Prompt-based memory: A-Mem (Xu et al., 2025) instantiates a memory system with construction, evolution, and retrieval agents; CAM (Li et al., 2025) and MemoryBank (Zhong et al., 2024) build multi-level (fine-to-coarse granularity) memory via multi-agent dedicated prompts.
-  • RL-based memory: Mem1 (Zhou et al., 2026) uses RL to optimize agents for memory searching during reasoning; Memory-R1 (Yan et al., 2025b) separately optimizes memory construction and retriever agents via RL.
-Figure 4: Detailed performance of different methods across seven question types on the PersonaMem benchmark across three history length settings (i.e., 32K, 128K, and 1M).
-The number of question types is consistent with those listed in Section 5.1
+Detailed information of baselines are provided in Appendix B.
+  • No memory: “Base” represents directly injecting raw conversations as LLMs’ input for query answering; RAG RAG answers queries by retrieving raw conversational history;
+  • Prompt-based memory: A-Mem xu2025amem instantiates a memory system with construction, evolution, and retrieval agents; CAM li2025cam and MemoryBank memorybank build multi-level (fine-to-coarse granularity) memory via multi-agent dedicated prompts.
+  • RL-based memory: Mem1 mem1 uses RL to optimize agents for memory searching during reasoning; Memory-R1 mem-r1 separately optimizes memory construction and retriever agents via RL.
 Implementation Details.
-To evaluate CoMAM’s effectiveness and generalizability, we conduct experiments on PersonaMem across its three history length settings (i.e., 32K, 128K, and 1M tokens) using LLM backbones from the Qwen (Bai et al., 2023) and Llama (Touvron et al., 2023) families.
+To evaluate CoMAM’s effectiveness and generalizability, we conduct experiments on PersonaMem and LongMemEval using LLM backbones from the Qwen qwen and Llama Llama families.
 Specifically, the extraction agent, profile agent, and reward model $\mathcal{V}$ are instantiated with Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct, while the retrieval agent leverages Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct, with model sizes tailored to each agent’s task complexity.
-All LLM backbones are configured with a 32K-token context window, and we adopt the average query-answer accuracy as the personalization evaluation metric.
+All LLM backbones are configured with a 32K-token context window.
+Following existing work personamem ; mem-r1 ; mem1 , we evaluate memory systems using query-answer accuracy.
+Specifically, we adopt multiple-choice accuracy for PersonaMem, and $F_{1}$, $B_{1}$, and $J_{1}$ (LLM-as-a-Judge) for LongMemEval.
 Table 1 presents the main results for both Qwen and Llama; other tables show only Qwen for simplicity.
@@
 5.2 Overall Performance (RQ1)
-To validate CoMAM’s efficacy for personalized LLMs, we compare its query-answer accuracy against leading memory system baselines on the PersonaMem benchmark.
-As shown in Table 1, CoMAM delivers superior performance across all context length configurations (32K, 128K, and 1M) and two LLM backbone families (Qwen and Llama), fully confirming CoMAM’s effectiveness and generalizability in personalized conversational scenarios.
+To validate CoMAM’s efficacy for the memory system, we compare its query-answer accuracy against leading memory system baselines on the PersonaMem benchmark across its three history length settings (i.e., 32K, 128K, and 1M tokens).
+As shown in Table 1, CoMAM delivers superior performance across all context length configurations (32K, 128K, and 1M) and two LLM backbone families (Qwen and Llama), fully confirming CoMAM’s effectiveness and generalizability in long-horizon conversational scenarios.
 Notably, RL-based memory baselines outperform prompt-based memory systems, validating that reinforcement learning is effective for optimizing memory system agents.
-For instance, the top prompt-based baseline CAM (Li et al., 2025) achieves only 0.450.45-$0.53$ accuracy across contexts, while the RL-based Memory-R1 (Yan et al., 2025b) maintains a higher accuracy range of 0.570.570.610.61.
+For instance, the top prompt-based baseline CAM li2025cam achieves only 0.450.45-$0.53$ accuracy across contexts, while the RL-based Memory-R1 mem-r1 maintains a higher accuracy range of 0.570.570.610.61.
 Moreover, CoMAM outperforms the state-of-the-art baseline Memory-R1, which employs independent agent optimization.
 This highlights the critical importance of joint optimization for multi-agent memory systems to achieve collaboration.
-Figure 4 details Qwen-based results for the seven question types, and we observe that CoMAM performs well in both detailed factual recall (e.g., Type 1) and abstract preference-based question types (e.g., Type 6), providing additional validation of its efficacy.
+Figure 6 in Appendix C.2 details Qwen-based results for the seven question types of PersonaMem.
+We further evaluate CoMAM on the LongMemEval benchmark.
+As shown in Table 2, our method achieves the best performance across all metrics, consistently outperforming strong baselines such as GAM GAM , further demonstrating its effectiveness in long-context memory retrieval and reasoning scenarios.
 Table 3: Ablation study for the contribution of different agents to the memory system.
@@
 w/o Re 0.54 0.51 0.51
-   mo train    0.50    0.46    0.44
+no train 0.50 0.46 0.44
 w/o train Ex 0.57 0.61 0.54
@@
 Ours 0.64 0.70 0.66
-Table 3: Ablation study of MDP regularization for heterogeneous, asynchronous agents. The “no train” variant denotes a memory system composed of agents merely guided by prompting.
+Table 4: Ablation study of joint optimization for heterogeneous agents.
 Method Accuracy (%)
@@
 + independent RL 0.57 0.56 0.54
-+ MDP RL 0.60 0.59 0.59
++ end-to-end RL 0.60 0.59 0.59
 Heterogeneous policy
@@
 + independent RL 0.57 0.61 0.59
-+ MDP RL (ours) 0.64 0.70 0.66
++ end-to-end RL (ours) 0.64 0.70 0.66
 5.3 Ablation for Agents (RQ2)
 To verify the necessity and contribution of the three core agents (Extraction Agent, Profile Agent, and Retrieval Agent), we designed targeted ablation experiments, with results presented in Table 4.
-First, the “Base” baseline (without memory agents) achieves the lowest performance across all context lengths, confirming the significance of memory systems for long-context personalized conversations.
+First, the “Base” baseline (without memory agents) achieves the lowest performance across all context lengths, confirming the significance of memory systems for long-horizon conversations.
 Second, removing any agent can lead to significant performance degradation, whether the Extraction Agent (“w/o Ex”) that builds factual fine-grained memories, the Profile Agent (“w/o Pr”) that captures core user preferences, or the Retrieval Agent (“w/o Re”) that searches for memories to answer queries.
-These results demonstrate that all three agents play complementary and unique roles in supporting the personalized memory system.
+These results demonstrate that all three agents play complementary and unique roles in supporting the memory system.
 Furthermore, the “w/o train” variants (full agent architecture with one agent left untrained) provide additional validation: untraining any single agent results in 2–8% performance drops compared to CoMAM, highlighting that optimizing these agents is critical to building an effective memory system.
-5.4 Ablation for MDP Regularization (RQ3)
-To address heterogeneity and asynchrony for joint optimization, we regularize agents’ execution as a sequential MDP trajectory.
-To validate MDP Regularization’s efficacy, we design “single policy” (all agents share one policy across tasks (Yu et al., 2026; Bo et al., 2024b)) and “heterogeneous policy” (each agent has a dedicated policy per task) variants, with two training strategies: “independent RL” (fine-tunes one agent’s policy at a time, others fixed without updates) and “MDP RL” (simultaneously updates policies via MDP trajectory for joint optimization).
+We further analyze the sensitivity of the reward coefficients $\alpha$ and $\beta$ in Figure 3, showing that the chosen settings ($\alpha=0.8$, $\beta=0.2$) enable the extraction agent to better preserve key information and the retrieval agent to effectively filter noise.
+Figure 3: Sensitivity of CoMAM to the hyperparameters $\alpha$ and $\beta$ as introduced in Sec. 4.1.
+5.4 Ablation for Dependency Modeling (RQ3)
+To model inter-agent dependencies for joint optimization, we formulate the heterogeneous multi-agent pipeline as a sequential MDP, which enables end-to-end RL.
+To validate its efficacy, we design “single policy” (all agents share one policy across tasks memagent ; share_one_policy ) and “heterogeneous policy” (each agent has a dedicated policy per task) variants, with two training strategies: “independent RL” (fine-tunes one agent’s policy at a time, others fixed without updates) and “end-to-end RL” (simultaneously updates policies via MDP trajectory for joint optimization).
 As shown in Table 4, “single policy” variants consistently underperform “heterogeneous policy” variants, confirming that a shared policy undermines agent specialization by integrating diverse task-specific objectives.
-Furthermore, “+MDP RL” outperforms “independent RL” across both policy paradigms, validating that MDP regularization explicitly models inter-agent sequential dependencies and enables both local task specialization and global collaboration.
+Furthermore, “+end-to-end RL” outperforms “independent RL” across both policy paradigms.
+This suggests that sequential MDP modeling during end-to-end RL training acts as a structured abstraction to capture how upstream decisions influence downstream behavior, thereby enabling more effective co-adaptation among agents.
 5.5 Ablation for Adaptive Credit Assignment (RQ4)
-To address RQ4, we design diverse credit assignment strategies to validate CoMAM’s adaptive mechanism.
-As shown in Table 4, optimizing agents solely with local task-specific rewards (“MDP RL w/ L”) or global rewards (“MDP RL w/ G”) yields performance improvements over the “no train” variant and “independent RL” variant, but lags behind variants that integrate global rewards with local rewards (“MDP RL w/ LG” and “Ours”).
-This demonstrates the necessity of combining local task rewards and global rewards to balance agent specialization on local tasks and collaboration for global performance.
-Notably, “Ours” (which is equipped with adaptive credit assignment) outperforms “MDP RL w/ LG” (which allocates equal shares of global rewards to local tasks), confirming the critical value of adaptively aligning local improvements with global performance.
-As illustrated in Figure 5, our adaptive weight $w_{n}$ for credit assignment outperforms any fixed weights, further validating the efficacy of the adaptive mechanism.
-Moreover, adaptive credit assignment can also enhance agents’ task-specific ability: our method outperforms both the “w/ LG” and “w/ L” on local task performance, validating this effect.
+To address RQ4, we evaluate different credit assignment strategies to assess the effectiveness of CoMAM’s adaptive mechanism.
+As shown in Table 6, optimizing agents using only local rewards (“end-to-end RL w/ L”) or only global rewards (“end-to-end RL w/ G”) improves over the “no train” and “independent RL” variants, but underperforms approaches that combine both signals (“end-to-end RL w/ LG” and “Ours”).
+This suggests that integrating local and global rewards is essential for balancing agent specialization and collaboration.
+Notably, “Ours”, with adaptive credit assignment, consistently outperforms “end-to-end RL w/ LG”, which uses uniform global rewards, highlighting the advantage of assigning credit based on agents’ contributions.
+As further shown in Figure 4, adaptive weights $w_{n}$ outperform fixed weighting schemes.
+Moreover, adaptive credit assignment improves task-specific performance: our method consistently achieves higher local task performance across all agents compared to both “w/ LG” and “w/ L”, indicating stronger role-specific capability.
+Additional experimental results and analyses of the reward curves are provided in Appendix C.1.
+5.6 Efficiency (RQ5)
+We further evaluate the training efficiency of CoMAM by comparing the convergence steps of joint optimization and independent optimization across 32K, 128K, and 1M context length settings, with results presented in Table 6.
+The total convergence steps serve as the core metric to reflect training efficiency, where fewer steps indicate higher efficiency in reaching stable performance.
+Specifically, independent optimization trains each agent separately, with total steps being the sum of individual convergence steps for the Extraction, Profile, and Retrieval Agents.
+In contrast, joint optimization trains all agents in parallel, with total steps determined by the slowest-converging agent, resulting in fewer total steps.
+While individual agents under joint training may take more steps for convergence (e.g., 58 vs.
+48 steps for the Extraction Agent at the 32K setting), joint training shortens the total training duration.
+Importantly, per-step cost is similar in both settings, since frozen agents in independent optimization still participate in the forward pass to complete the pipeline and compute rewards.
+Thus, the efficiency gain mainly comes from reducing redundant forward passes, as agents update simultaneously in joint optimization.
+Moreover, inference latency is decoupled from the training paradigm: agents optimized by joint or independent optimization achieve comparable inference efficiency at deployment.
+Yet joint optimization delivers superior global performance via inter-agent collaboration and adaptive credit assignment, highlighting the distinct advantage of joint multi-agent optimization for memory systems.
+Figure 4: The left demonstrates CoMAM’s sensitivity to the credit assignment weight, and the right shows the credit assignment’s impact on each local agent’s performance on the 32K setting.
 Table 5: Ablation study for different credit assignment strategies.
@@
 + independent RL 0.57 0.61 0.59
-MDP RL w/ L 0.57 0.65 0.61
-MDP RL w/ G 0.60 0.64 0.60
-MDP RL w/ LG 0.62 0.65 0.63
+end-to-end RL w/ L 0.57 0.65 0.61
+end-to-end RL w/ G 0.60 0.64 0.60
+end-to-end RL w/ LG 0.62 0.65 0.63
 Ours 0.64 0.70 0.66
-5.6 Efficiency (RQ5)
-We further evaluate the training efficiency of CoMAM by comparing the convergence steps of joint optimization and independent optimization across 32K, 128K, and 1M context length settings, with results presented in Table 5.
-The total convergence steps serve as the core metric to reflect training efficiency, where fewer steps indicate higher efficiency in reaching stable performance.
-Specifically, independent optimization trains each agent separately, with total steps being the sum of individual convergence steps for the Extraction, Profile, and Retrieval Agents.
-In contrast, joint optimization trains all agents in parallel, with total steps determined by the slowest-converging agent, resulting in fewer total steps.
-While individual agents under joint training may take more steps (e.g., 58 vs.
-48 steps for the Extraction Agent at the 32K setting), parallel training shortens the total training duration.
-Notably, inference latency is decoupled from the training paradigm: agents optimized by joint or independent optimization achieve comparable inference efficiency at deployment.
-Yet joint optimization delivers superior global performance via inter-agent collaboration, highlighting the distinct advantage of joint multi-agent optimization for personalized memory systems.
-Figure 5: The left demonstrates CoMAM’s sensitivity to the credit assignment weight, and the right shows the credit assignment’s impact on each local agent’s performance on the 32K setting.
 Table 6: Convergence step comparison of the joint and independent optimization.
@@
 6 Conclusion
-In this paper, we propose CoMAM, a collaborative reinforcement learning framework for jointly optimizing agents in personalized memory systems.
-It tackles the issue that independent agent optimization can not ensure global system performance without cross-agent collaboration.
-To enable joint optimization, we regularize the asynchronous executions of heterogeneous agents as a sequential MDP, which embeds inter-agent dependencies into state transitions and yields both local and global rewards.
-To align local improvement with global performance, we introduce an adaptive credit assignment mechanism that allocates global rewards to individual agents based on group-level ranking consistency between global and local rewards.
-Experiments on long-term personalized conversational benchmarks demonstrate that CoMAM outperforms leading memory baselines, validating the effectiveness of our collaborative RL framework.
+In this work, we study the limitations of independently optimizing agents in multi-agent memory systems, highlighting the lack of co-adaptation and the ambiguity of credit assignment under a shared global objective.
+To address these challenges, we propose CoMAM, a joint optimization framework that enables agents’ co-adaptation via end-to-end reinforcement learning, together with an adaptive credit assignment mechanism for more informative training signals.
+By modeling the inter-agent dependencies as an MDP and integrating local and global rewards adaptively, CoMAM facilitates collaboration across agents.
+Experimental results demonstrate consistent improvements over strong baselines, validating the effectiveness of joint optimization with adaptive credit assignment for multi-agent memory systems.
 The limitation is discussed in Appendix E.
-Impact Statements
-This paper presents work whose goal is to advance the field of machine learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Algorithm
@@
 8:  $r_{\text{ret}}\leftarrow\beta\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}|}+(1-\beta)\cdot\frac{|\mathcal{E}_{q}^{\prime}\cap\mathcal{E}_{q}|}{|\mathcal{E}_{q}^{\prime}\cup\mathcal{E}_{q}|}$
-9:  $r_{\text{ans}}\leftarrow\mathbb{I}(p^{*}=p)$ $\triangleright$ Global reward (accuracy)
+9:  $r_{\text{ans}}\leftarrow\mathbb{I}(p^{*}=p)$ $\triangleright$ Global reward (Downstream task accuracy)
 10:end for
@@
 Each session contains 15–30 conversation turns, designed to naturally reveal or update user preferences.
-In addition, each history is paired with a suite of subsequent queries across seven types, enabling comprehensive evaluation of personalized response capabilities.
+In addition, each history is paired with a suite of subsequent queries across seven types, enabling comprehensive evaluation of the memory system.
+The question types include: (1) Recall user-shared facts, (2) Suggest new ideas, (3) Acknowledge latest user preferences, (4) Track full preference evolution, (5) Revisit reasons for preference updates, (6) Provide preference-aligned recommendations, and (7) Generalize to new scenarios.
+Compared to existing benchmarks [19, 32], PersonaMem’s strengths in long context history and different query type enable rigorous assessment of memory systems in conversational scenarios.
 For PersonaMem’s three context length settings (32k, 128k, and 1M tokens), we partition their training and validation sets based on these queries.
 The statistical breakdown of training and validation queries is summarized in Table 7.
+In the LongMemEval benchmark [32], each sample consists of a scalable multi-session user–assistant interaction history, where information is gradually revealed and may be revised across sessions.
+The benchmark contains 500 manually created questions and defines two standard settings: LongMemEval-S with about 115K tokens per sample, and LongMemEval-M with 500 sessions and roughly 1.5M tokens.
+Its evaluation targets five core memory abilities—information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention—and the released question categories include single-session-user, single-session-preference, single-session-assistant, multi-session, temporal-reasoning, and knowledge-update.
+Compared with benchmarks that mainly test long-context recall, LongMemEval more explicitly evaluates cross-session reasoning and dynamic user-state tracking in realistic assistant settings.
 To process the training set, we first identify the specific session within the full conversational history that contains explicit evidence relevant to each query $q$, and we define this session as the query’s targeted conversational history $\mathcal{H}$.
@@
 Ground-truth evidence $\mathcal{E}_{q}$ for each query $q$ is extracted using the Seed-1.6 model and validated through human evaluation.
+We emphasize that $\mathcal{E}_{q}$ is only used for computing training rewards and is not available during inference, ensuring no information leakage.
+While obtaining $\mathcal{E}_{q}$ involves additional annotation effort, this cost is incurred only during training and does not affect inference-time efficiency, making it a practical and acceptable design choice.
 For evaluation, multiple distinct queries may share the same full conversational history that includes multiple sessions.
@@
 Specifically, when the length of the input conversation exceeds the context window constraint, the input is truncated at the context window limit.
-  • RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history via embedding-based similarity matching.
+  • RAG [14] answers queries by retrieving raw conversational history via embedding-based similarity matching.
 The embedding model employed is all-MiniLM-L6-v2, and ultra-long input contexts are uniformly partitioned into fixed-length segments, each consisting of 2,048 tokens.
 Prompt-based memory systems
-  • A-Mem (Xu et al., 2025) instantiates a memory system through multiple agents with dedicated prompts, including the memory retrieval construction, precursor memory evolution, and retrieval agents;
-  • CAM (Li et al., 2025) builds a hierarchical memory system via multiple agents, where dedicated agents undertake node construction (via ego-centric disentanglement and node replication) and information aggregation (for hierarchical abstraction).
+  • A-Mem [35] instantiates a memory system through multiple agents with dedicated prompts, including the memory retrieval construction, precursor memory evolution, and retrieval agents;
+  • CAM [15] builds a hierarchical memory system via multiple agents, where dedicated agents undertake node construction (via ego-centric disentanglement and node replication) and information aggregation (for hierarchical abstraction).
 We implement it on the PersonaMem benchmark.
-  • MemoryBank (Zhong et al., 2024) construct a human-like memory for LLMs that uses a forgetting curve to remember and forget information over time.
-Similarly, we implement it on the PersonaMem benchmark.
+  • MemoryBank [42] construct a human-like memory for LLMs that uses a forgetting curve to remember and forget information over time. Similarly, we implement it on the PersonaMem benchmark.
+  • SeCom [21] designs a modular memory system for long-horizon conversational agents, where dedicated components handle memory construction and retrieval using prompt-based multi-agent.
+  • Mem0 [6] instantiates a multi-agent memory system that maintains fine-grained and coarse-grained user memories, performing extraction and retrieval with specialized prompts.
 RL-based memory:
-  • Mem1 (Zhou et al., 2026) proposes a reinforcement learning (RL) framework that empowers agents to conduct efficient multi-turn information retrieval and reasoning for query answering.
+  • GAM [37] introduces an agentic memory system with a memorizer for key information and a researcher for retrieval and integration at runtime, trained with reinforcement learning.
+  • Mem1 [44] proposes a reinforcement learning (RL) framework that empowers agents to conduct efficient multi-turn information retrieval and reasoning for query answering.
 The optimized agents prioritize direct memory retrieval from history, rather than memory construction.
-  • Memory-R1 (Yan et al., 2025b) constructs a Memory Manager that learns optimal memory operations (ADD, UPDATE, DELETE) and an Answer Agent tasked with distilling relevant information from retrieved entries.
+  • Memory-R1 [38] constructs a Memory Manager that learns optimal memory operations (ADD, UPDATE, DELETE) and an Answer Agent tasked with distilling relevant information from retrieved entries.
 Uniquely, the system is optimized via independent reinforcement learning (RL) in two stages: the Memory Manager is optimized first, followed by the Answer Agent.
@@
 For a fair comparison, all the baselines are implemented with the same LLM backbones as our CoMAM, including Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct.
+All experiments are conducted on a cluster of 8 NVIDIA A100 GPUs.
 For the implementation of variant “+independent RL”, each agent is trained by fusing its task-specific rewards and global accuracy rewards, with the remaining agents frozen.
 For the “single policy” variant, we adopt a shared policy to serve as the policy for all three agents during training and inference.
-The prompts for different agents are presented below.
+The prompts for specializing different agents and rubrics for evaluating the profile agent are presented below.
 The temperature of LLMs is set to $1$ for training and $0.8$ for evaluation.
@@
 Please output the retrieved relevant information in <information> </information> and choose the most appropriate option (a), (b), (c), or (d) in <final_answer></final_answer>.
+Rubrics for Profile Agent Reward Evaluation You are an evaluation expert.
+Your task is to assess whether the provided Coarse-grained memory accurately captures the key information from the given Conversation History.
+Conversation History:
+  User--LLM conversational history
+Coarse-grained memory:
+  Output generated by the Profile Agent
+Evaluation Criteria: 1.
+Are the scenarios sufficiently rich and comprehensive?
+2.
+Is the analysis of user preferences sufficiently in-depth?
+3.
+Is the user persona sufficiently detailed and representative?
+Output Format: Please output a single scalar score within <score></score> as a float between 0.0 and 1.0, without any additional text.
+1.0: The synthesized insight perfectly captures the scenario, user preferences, and persona.
+0.0: The synthesized insight fails to capture the relevant information.
 Appendix C More Detailed Experimental Results
+Figure 5: The curves for local task-specific reward $r_{n}$ and global accuracy reward $r_{3}$.
 C.1 The Impact of Joint Optimization on Local Task Specialization
@@
 This indicates that adaptive credit assignment will not compromise local task capabilities; instead, it strengthens them by aligning local improvement with global goals.
-  • Global-local Collaboration: The global reward curve of “Ours” converges to 0.70, significantly higher than “w/ L’ (0.65) and “w/ G” (0.64), proving that joint optimization realizes global performance enhancement by fostering inter-agent collaboration.
+  • Global performance enhancement: The global reward curve of “Ours” converges to 0.70, significantly higher than “w/ L’ (0.65) and “w/ G” (0.64), proving that joint optimization realizes global performance enhancement by fostering inter-agent collaboration.
 Adaptive credit assignment mechanism quantifies the correlation between each agent’s local performance and the global outcome, ensuring that agents receive rewards not only for completing local tasks well but also for contributing to the global goal.
-This guides agents to optimize their local strategies in a way that benefits the entire system, thus achieving both local specialization and global collaboration.
-Figure 6: The curves for local task-specific reward $r_{n}$ and global accuracy reward $r_{3}$.
+This guides agents to optimize their local strategies in a way that benefits the entire system, thus achieving both local specialization and inter-agent collaboration for the global memory system.
+C.2 Detailed Performance Across Question Types
+Figure 6 details Qwen-based results for the seven question types, and we observe that CoMAM performs well in both detailed factual recall (e.g., Type 1) and abstract preference-based question types (e.g., Type 6), providing additional validation of its efficacy.
+Figure 6: Detailed performance of different methods across seven question types on the PersonaMem benchmark across three history length settings (i.e., 32K, 128K, and 1M).
+The number of question types is consistent with those listed in Section 5.1
 Appendix D Discussion
@@
     Existing frameworks primarily adopt independent optimization.
-For example, MIRIX (Wang & Chen, 2025) designs a meta-agent and six specialized agents for memory management, optimizing each via tailored prompts; Memagent (Yu et al., 2026) focuses on a single retrieval agent, training it with query-answer accuracy as the reward; Memory-R1 (Yan et al., 2025b) splits the system into a Memory Manager and Answer Agent, training them in separate stages—first optimizing the Manager with the Agent frozen, then freezing the Manager to optimize the Agent.
-This approach ignores inter-agent collaboration: memory construction is not adjusted based on retrieval needs, and the Agent can only adapt to a fixed Manager’s output, leading to “local optimal but global suboptimal” outcomes.
-    CoMAM proposes end-to-end joint optimization via MDP regularization.
-We model the asynchronous execution of Extraction, Profile, and Retrieval Agents as a sequential MDP trajectory, where each agent’s output directly serves as the next agent’s input.
-All three agents are updated simultaneously during training, with inter-agent dependencies embedded into state transitions, ensuring they evolve collaboratively toward the global goal of query-answer accuracy.
+For example, MIRIX [29] designs a meta-agent and six specialized agents for memory management, optimizing each via tailored prompts; Memagent [39] focuses on a single retrieval agent, training it with query-answer accuracy as the reward; Memory-R1 [38] splits the system into a Memory Manager and Answer Agent, training them in separate stages—first optimizing the Manager with the Agent frozen, then freezing the Manager to optimize the Agent.
+This approach ignores inter-agent independencies and lacks co-adaptation: memory construction is not adjusted based on together with retrieval behavior, and the agent can only adapt to a fixed Manager’s output, leading to insufficient inter-agent collaboration.
+    CoMAM proposes end-to-end joint optimization via modeling inter-agent dependencies in the MDP trajectory, where each agent’s output directly serves as the next agent’s input.
+All three agents are updated simultaneously during training, with inter-agent dependencies embedded into state transitions, ensuring they evolve collaboratively toward the global goal of the memory system.
   • Reward Design: Single Global Reward vs. Adaptive Credit Assignment
     Existing frameworks that optimize memory agents with RL primarily rely solely on final query-answer accuracy as the reward for all agents.
-For example, Mem-$\alpha$ and Mem1 (Zhou et al., 2026) use this single reward to train construction and retrieval agents respectively.
-Memory-R1 (Yan et al., 2025b) and GAM (Yan et al., 2025a) also adopt this design, using only answer accuracy to optimize their agents separately.
+For example, Mem-$\alpha$ and Mem1 [44] use this single reward to train construction and retrieval agents, respectively.
+Memory-R1 [38], and GAM [37] also adopt this design, using only answer accuracy to optimize their agents separately.
 This ambiguous credit assignment prevents agents from distinguishing their individual contributions, leading to unstable convergence.
@@
 Homogeneous Policy: Some multi-agent reinforcement learning (RL) methods adopt a shared policy across all agents.
-For instance, CORY (Ma et al., 2024) proposes an RL framework featuring “Pioneer” and “Observer” agents that share a single policy model.
+For instance, CORY [18] proposes an RL framework featuring “Pioneer” and “Observer” agents that share a single policy model.
 These agents co-evolve through multi-agent collaboration and periodic role-swapping.
-In contrast, considering the heterogeneity and asynchrony of a memory system, CoMAM employs heterogeneous policies tailored to distinct agents.
+In contrast, considering the heterogeneity of agents in a memory system, CoMAM employs heterogeneous policies tailored to distinct agents.
   • Workflow: Sequential vs.
 Concurrent: Existing multi-agent system (MAS) methods often adopt a debate-style pipeline for reasoning tasks, such as math tasks.
-For instance, MAPoRL (Park et al., 2025) models agent interactions as concurrent debate, where multiple agents generate responses in parallel, focusing on persuasive collaboration for general reasoning.
-In contrast, CoMAM explicitly formulates agents’ executions as a sequential Markov Decision Process (MDP), embedding inter-agent dependencies into state transitions, ensuring that memory construction quality dynamically adapts to retrieval requirements and vice versa.
+For instance, MAPoRL [22] models agent interactions as concurrent debate, where multiple agents generate responses in parallel, focusing on persuasive collaboration for general reasoning.
+In contrast, CoMAM explicitly models inter-agent dependencies as a sequential Markov Decision Process (MDP), providing a structured way to expose how upstream decisions influence downstream behavior.
+This ensures that memory construction quality dynamically adapts to retrieval requirements and vice versa.
   • Collaboration Paradigm: Joint Optimization vs.
-Independent Alignment: Existing methods, such as Optimas (Wu et al., 2026), foster collaboration by aligning local rewards with global objectives via heuristic reward shaping for compound AI systems.
+Independent Alignment: Existing methods, such as Optimas [33], foster collaboration by aligning local rewards with global objectives via heuristic reward shaping for compound AI systems.
 However, it optimizes individual agents separately, leveraging locally-global aligned rewards rather than through end-to-end joint training.
-In contrast, CoMAM adopts a joint optimization paradigm: it regularizes the asynchronous execution of heterogeneous agents into a sequential MDP trajectory and updates all agents simultaneously.
+In contrast, CoMAM adopts a joint optimization paradigm that updates all agents simultaneously.
   • Credit Assignment: Adaptive Assignment vs.
-Equal Allocation: Most existing systems use static credit assignment mechanisms without quantifying agents’ varying contributions: AT-GRPO (Zhao et al., 2026) distributes rewards equally across agents using fixed weights, ignoring task differences; CoMAS (Xue et al., 2026) adopts a zero-sum adversarial reward scheme that pits agents against each other, which is misaligned with the cooperative nature of memory systems; In contrast, CoMAM’s credit assignment is grounded in group-level ranking consistency (NDCG) between each agent’s local task rewards (e.g., coverage for extraction, rationality for profiling, precision for retrieval) and the global system reward (query-answer accuracy).
-This design quantifies each agent’s contribution to the global system.
-It dynamically adjusts weights according to agent performance, ensuring adaptive credit allocation that aligns local improvements with global performance.
+Equal Allocation: Most existing systems use static credit assignment mechanisms without quantifying agents’ varying contributions: AT-GRPO [40] distributes rewards equally across agents using fixed weights, ignoring task differences; CoMAS [36] adopts a zero-sum adversarial reward scheme that pits agents against each other, which is misaligned with the cooperative nature of memory systems; In contrast, CoMAM assigns agent-specific credit based on a ranking-based proxy that measures the alignment between each agent’s local task rewards and the global system reward.
+This formulation captures whether trajectories with higher global performance are consistently associated with higher local rewards for a given agent.
+Agents with stronger alignment receive higher credit, as their behaviors are more predictive of overall system improvement.
+This provides more informative training signals for specialized agents.
 Appendix E Limitation
-While CoMAM enables effective joint optimization for multi-agent memory systems, it is worth noting that our current implementation focuses on verifying the core logic of collaborative RL, with certain extended capabilities reserved for future exploration to better adapt to more complex scenarios.
-1. Limited Agent Task Scope:
-The framework currently leverages a compact agent set (Extraction, Profile, Retrieval) to focus on verifying the efficacy of collaborative RL for personalized memory tasks.
-To prioritize the core verification goal, we reserve extended agent capabilities (such as memory modification for consistency maintenance and redundant memory deletion) for future work.
-2. Static MDP Regularization for Training:
-Current MDP regularization models agent execution as a static sequential trajectory, enabling joint training of heterogeneous and asynchronous agents.
-Compared with the existing independent optimization training paradigm, this design can better enhance the collaborative capabilities of agents beyond task-specific capabilities.
-While effective in boosting inter-agent collaboration, we can further consider the dynamic deployment of agents in real-world scenarios, such as dynamic invocation frequencies and real-time memory updates.
-Future work will integrate these real-world dynamics into the joint training process to improve the consistency between training and deployment.
-3. Lack of Explicit local-local Alignment:
-The adaptive credit assignment is specifically designed to align local agent improvements with global system goals, enabling agents’ collaboration towards the global goal.
-Since all agents’ local subtasks are oriented towards the consistent global system goal, the alignment between their local tasks is implicitly modeled.
-In the future, we will explore explicitly modeling this local-local alignment for agents’ local subtasks, which will further boost inter-agent collaboration.
+While CoMAM demonstrates the effectiveness of joint optimization with adaptive credit assignment for multi-agent memory systems, several limitations remain.
+1.
+Simplified Agent Design.
+We adopt a compact set of agents (Extraction, Profile, and Retrieval) to focus on core collaborative behaviors.
+More advanced functionalities, such as memory editing, consistency maintenance, and redundancy removal, are not explicitly modeled and remain for future exploration.
+2.
+Limited Cross-Agent Credit Modeling.
+CoMAM estimates agent-specific credit by modeling the ranking consistency between each agent’s local rewards and the global system reward.
+However, relationships among agents’ local rewards are not explicitly considered in the credit estimation.
+Incorporating such cross-agent relationships may provide more informative credit signals and further improve coordination.
+Appendix F Broader Impact
+CoMAM aims to improve the effectiveness of multi-agent memory systems for long-horizon conversational tasks by introducing a joint optimization framework with end-to-end reinforcement learning and adaptive credit assignment.
+This may benefit applications such as personalized assistants and long-context reasoning systems.
+However, it does not inherently address potential privacy risks associated with sensitive user information.